### v1

# Experimental Design for Active Transductive Inference in Large Language Models

Abstract
Transduction, the ability to include query-specific examples in the prompt at inference time, is one of the emergent abilities of large language models (LLMs). In this work, we propose a framework for adaptive prompt design called active transductive inference (ATI). We design the LLM prompt by adaptively choosing few-shot examples for a given inference query. The examples are initially unlabeled and we query the user to label the most informative ones, which maximally reduces the uncertainty in the LLM prediction. We propose two algorithms, GO and SAL, which differ in how the few-shot examples are chosen. We analyze these algorithms in linear models: first GO and then use its equivalence with SAL. We experiment with many different tasks and show that GO and SAL outperform other methods for choosing few-shot examples in the LLM prompt at inference time.
1 Introduction
Large language models (LLMs), such as Vicuna (Chiang et al., 2023), Falcon-40B (Penedo et al., 2023), and OpenLLaMA (Touvron et al., 2023) are applied in mainly two ways: fine-tuning and prompt tuning. In fine-tuning, the LLM weights are adapted to a downstream task (Devlin et al., 2018). fine-tuning allows for the inclusion of domain-specific knowledge that a pre-trained model may not possess and resembles classic inductive inference. Fine-tuned models often do not need carefully designed prompts, which makes them easier to deploy. The main drawback of fine-tuning is that it can be costly, because tens of thousands of training examples may be needed to fine-tune billions of parameters of the LLM (Ding et al., 2023). In prompt tuning, the LLM weights are fixed and the LLM is given query-specific examples at inference time that affect its output (Lester et al., 2021). This ability to conduct transductive inference is one of the emergent abilities of LLMs. Prompt tuning does not require large training sets. It is also preferred when query-specific examples are private or change over time, and thus can only be utilized at inference time. Prior works on prompt tuning mainly focus on hard prompt tuning, where the prompt is carefully handcrafted to get the desired output. This can be time-consuming, as even a simple change in the prompt can lead to poor performance on the downstream task (Suzgun et al., 2022). We propose a framework for adaptive prompt design called active transductive inference (ATI). The key idea is to design the LLM prompt by adaptively choosing few-shot examples for a given inference query. The examples are initially unlabeled and we query the user to label the most informative ones, which maximally reduce the uncertainty in the LLM prediction. To illustrate our approach, we show two exemplary scenarios next.
The first example is an LLM as a medical diagnostics chatbot. The LLM is tasked with identifying potential medical conditions based on symptoms. The user inputs that they experience fatigue, cough, and shortness of breath. These symptoms can be caused by various conditions: a common cold, flu, or something more severe like heart issues. Instead of immediately offering a range of potential remedies, the LLM could actively query the user with symptoms that are specific to a particular condition, such as chest pain, leg swelling, and irregular pulse for heart-related issues. It then asks the user to label these symptoms as present or absent. This interaction allows the LLM to refine its diagnosis.
[…]
Figure 1: A general workflow of active transductive inference (ATI). The active learning module selects examples sequentially to be labeled by a human. Since LLMs can be viewed as general inference machines, our approach is broadly applicable to classification, regression, general pattern recognition, and natural language tasks. The general pattern recognition tasks include Abstract Reasoning Task (ARC) (Alford, 2021) and identifying Probabilistic Context Free Grammar tasks (PCFG) (Hupkes et al., 2020).
To address the above challenges, we propose a framework for adaptive prompt design called active transductive inference (ATI). Our solution is general and can be easily extended to any active supervised-learning task, like active regression (Gao & Koller, 2011) and active classification (Gao & Koller, 2011) (see Figure 1). At a high level, we treat the LLM as a general inference machine that is shown adaptively-chosen examples with labels at inference time. The LLM then utilizes them to answer the original user’s query. The key idea is to choose the next example to label as the one that maximally reduces the estimated uncertainty of the answer to the user’s query. We focus on designing algorithms with the following two properties: (1) Implementable in any LLM that can be queried efficiently. The parameters of the LLM do not change or have to be observed. (2) Analyzable in simple models. In this work, we use linear models to motivate and analyze our algorithms.
We now state the main contributions of our work:
(1) We propose a G-Optimal design algorithm (GO). The key idea in GO is to query the user to label the closest examples to the example in the inference task. Our main contribution is the right notion of closeness, based on posterior covariance in a simpler model. GO is implementable with any LLM that can be sampled from, and does not require access to model parameters, feature embeddings of the LLM, or its gradients.
(2) We propose a Simulation-Based Active Learning algorithm (SAL). SAL uses simulation to estimate the impact of labeling unlabeled examples on uncertainty of the example in the inference task. SAL is also implementable with any LLM that can be sampled from.
(3) GO is motivated by optimal designs in linear models (Kiefer & Wolfowitz, 1960; Pukelsheim, 2006). This allows us to analyze GO in linear models (Theorem 4). Our proof is a major departure from similar analyses in fixed-budget best-arm identification in bandits (Azizi et al., 2022; Yang & Tan, 2022), for instance because we analyze the discrete allocation problem and each unlabeled point can be labeled at most once. We discuss this in detail right after Theorem 4. SAL is more general than GO because it does not make any linear model assumption in its design. We show that SAL and GO are equivalent in linear models in Theorem 5.
(4) We evaluate GO and SAL on UCI (Markelle Kelly, 1988) and OpenML (Vanschoren et al., 2013) regression and classification tasks, custom NLP datasets, abstract reasoning corpus (ARC) tasks (Alford, 2021; Mirchandani et al., 2023), and Probabilistic Context Free Grammar (PCFG) tasks (Hupkes et al., 2020). GO and SAL consistently outperforms other methods for choosing few-shot examples.
We advance the understanding of active transductive inference in LLMs and develop a practical methodology for adaptive prompt design. This paper is organized as follows. Section 2 introduces the problem setting. Section 3 presents our methods and discusses their properties. Section 4 is devoted to analyzing our methods. Section 5 validates our approach empirically. We review related work in detail in Appendix A. Finally, Section 6 summarizes our contributions and suggests avenues for future work.
2 Setting
We pose the problem of adaptive prompt design as active learning. We adopt the following standard active learning terminology (Dasgupta, 2005). We have $n$ unlabeled points $\mathcal{X}=\left\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\right\}\subset\mathbb{R}^{d}$. Each point is a vector in $d$ dimensions. The space of labels is $\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$. Each label is a vector in $d_{y}$ dimensions. Each point $\mathbf{x}\in\mathcal{X}$ has a label $Y\in\mathcal{Y}$. The labels are unknown. The point $\mathbf{x}$ and its label $Y$ are related through an unknown model parameter $\bm{\theta}_{*}$. Specifically, $Y=f(\mathbf{x},\bm{\theta}_{*})+\varepsilon$, where $f$ represents the model and $\varepsilon$ is independent zero-mean noise. The stochasticity of $Y$ allows us to model noisy labels even if $\bm{\theta}_{*}$ was known. One point in $\mathcal{X}$ is designated as a target point $\mathbf{x}_{*}$ and its target label is $Y_{*}=f(\mathbf{x}_{*},\bm{\theta}_{*})+\varepsilon_{*}$. We want to infer it by asking a human to label other points in $\mathcal{X}$.
We illustrate our notation next. In theme recognition, $\mathbf{x}$ is a set of similar items, $Y$ is their theme, and $\bm{\theta}_{*}$ is a model parameter that relates $\mathbf{x}$ and $Y$. Since the model parameter $\bm{\theta}_{*}$ is unknown, the target label $Y_{*}$ is unknown. To reduce its uncertainty, and ultimately learn it, we actively query the user for labels of carefully chosen points $\mathbf{x}\in\mathcal{X}$. We reduce the uncertainty of $Y_{*}$ through that of $\bm{\theta}_{*}$. To understand how we achieve it, we need additional notation. We related the points and their labels, conditioned on previously labeled points, using large language models. Specifically, let $H_{t}=(X_{\ell},Y_{\ell})_{\ell\in[t-1]}$ be the history of $t-1$ previously labeled points, where $X_{\ell}$ is the $\ell$-th labeled point and $Y_{\ell}$ is its label. Let $p(\cdot\mid\mathbf{x},H_{t})$ be the distribution of the label of point $\mathbf{x}$ conditioned on $H_{t}$. Sampling from this distribution is implemented implicitly by the LLM, where $H_{t}$ is provided as $t-1$ few-shot examples in the LLM prompt. Then it is queried for the label of $\mathbf{x}$. We visualize this in Figure 1
A useful way of thinking of $p(\cdot\mid\mathbf{x},H_{t})$ is as a Bayesian posterior distribution (Xie et al., 2021; Deng et al., 2023; Lee et al., 2023) associated with a graphical model $\mathbf{x}\to Y\leftarrow\bm{\theta}_{*}\to H_{t}$. This distribution can be sampled from as follows. First, $\bm{\theta}_{*}$ is sampled conditioned on $H_{t}$, which accounts for model uncertainty. Then $Y$ is sampled conditioned on $\bm{\theta}_{*}$ and $\mathbf{x}$, which accounts for additional label uncertainty. In this model, reducing uncertainty in $\bm{\theta}_{*}$ makes $Y_{*}$ more certain. In LLMs, the model parameter $\bm{\theta}_{*}$ is typically not exposed but sampling $Y\sim p(\cdot\mid\mathbf{x},H_{t})$ can be implemented efficiently. The prior distribution of labels is $p(\cdot\mid\mathbf{x},H_{1})$ where $H_{1}=\left\{\right\}$.
To reduce the uncertainty of $Y_{*}\mid\mathbf{x}_{*},H_{t}$, we need to measure it. One option is $-\mathbb{E}_{Y_{*}\sim p(\cdot\mid\mathbf{x}_{*},H_{t})}[\log p(Y_{*}\mid\mathbf{x}_{*},H_{t})]$. The main issue with entropy is that it is hard to estimate for high-dimensional random variables (Vershynin, 2020), especially without having access to $p(\cdot\mid\mathbf{x}_{*},H_{t})$ beyond sampling from it. Therefore, we opt for $\mathrm{cov}[Y_{*}\mid\mathbf{x}_{*},H_{t}]$, which can be estimated using as few as two samples from $p(\cdot\mid\mathbf{x}_{*},H_{t})$. Specifically, we measure the uncertainty of $Y_{*}\mid\mathbf{x}_{*},H_{t}$ by $\operatorname{tr}(\mathrm{cov}[Y_{*}\mid\mathbf{x}_{*},H_{t}])$ (Pukelsheim, 2006). Note that this is the sum of variances of all entries of $Y_{*}\mid\mathbf{x}_{*},H_{t}$. For scalar $Y_{*}$, this becomes $\mathrm{var}[Y_{*}\mid\mathbf{x}_{*},H_{t}]$, a natural measure of uncertainty in linear regression.
One naive approach to reducing the uncertainty of $Y_{*}$ is to label all points in $\mathcal{X}$ and include them in the LLM prompt. This has two issues. First, it is statistically inefficient because all points in $\mathcal{X}$ must be labeled. Second, since LLM prompts have a limited length, not all labeled points may fit. Therefore, we study how to intelligently choose a small subset of $\mathcal{X}$ of size $T$ that would fit in the prompt.
Our active transductive inference (ATI) problem is formally defined as follows. Fix a budget $T$. We then design a sequential adaptive algorithm over $T$ rounds, where the point $X_{t}$ in round $t\in[T]$ is chosen as a function of history $H_{t}$ up to that round. The label of point $X_{t}$ is $Y_{t}=f(X_{t},\bm{\theta}_{*})+\varepsilon_{t}$, where $\varepsilon_{t}$ is independent zero-mean noise in round $t$, and $Y_{t}$ is initially unknown. This makes our setting strictly more general than choosing the best set of labeled examples for prompt design (Lester et al., 2021; Dong et al., 2022; Zhang et al., 2022a; Liu et al., 2023a; Yu et al., 2023; Liu et al., 2023b; Diao et al., 2023). Our goal is to minimize the uncertainty $\operatorname{tr}(\mathrm{cov}[Y_{*}\mid\mathbf{x}_{*},H_{T+1}])$.
3 Algorithms
This section has four subsections. In Sections 3.1 and 3.2, we introduce our active learning algorithms. The algorithms are proposed for a single target point with a scalar label. In Section 3.3, we generalize the algorithms to vector labels. In Section 3.4, we extend them to multiple target points.
3.1 Optimal Design Algorithm
Algorithm 1 G-optimal design (GO)
1:  Input: Target point $\mathbf{x}_{*}$, budget $T$
2:
3:  $\mathcal{L}_{1}\leftarrow\emptyset,\ \mathcal{U}_{1}\leftarrow[n],\ H_{1}\leftarrow()$
4:  for $t=1,\dots,T$ do
5:     $I_{t}\leftarrow\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{*}$
6:     $X_{t}\leftarrow\mathbf{x}_{I_{t}}$
7:     Observe label $Y_{t}$ of point $X_{t}$
8:     $\mathcal{L}_{t+1}\leftarrow\mathcal{L}_{t}\cup\{I_{t}\},\ \mathcal{U}_{t+1}\leftarrow\mathcal{U}_{t}\setminus\{I_{t}\}$
9:     $H_{t+1}\leftarrow H_{t}\oplus(X_{t},Y_{t})$
10:  end for
11:
12:  Output: Sample $Y_{*}\sim p(\cdot\mid\mathbf{x}_{*},H_{T+1})$
The key idea is to label similar points $\mathbf{x}$ to the target point $\mathbf{x}_{*}$. The notion of similarity is derived based on insights from a linear model. Specifically, let $Y\sim\mathcal{N}(\mathbf{x}^{\top}\bm{\theta}_{*},\sigma^{2})$ and $\bm{\theta}_{\ast}\sim\mathcal{N}(\bm{\theta}_{0},\mathbf{\Sigma}_{0})$, where $\bm{\theta}_{0}$ and $\mathbf{\Sigma}_{0}$ are the prior mean and covariance of the unknown model parameter $\bm{\theta}_{*}$, respectively. Then for any history $H_{t}$, the posterior variance of $\mathbb{E}[Y_{*}\mid\mathbf{x}_{*},H_{t}]$ is $\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t}\mathbf{x}_{*}$, where
$\widehat{\mathbf{\Sigma}}_{t}=\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}\right)^{-1}$
is the posterior covariance of $\bm{\theta}_{*}\mid H_{t}$. This is a classic result for Bayesian inference in linear models (Bishop, 2006).
One approach to minimize the uncertainty of $Y_{*}\mid\mathbf{x}_{*},H_{t}$ is to label points that minimize it greedily. Let $\mathcal{L}_{t}\subseteq[n]$ and $\mathcal{U}_{t}\subseteq[n]$ be the indices of labeled and unlabeled points up to round $t$, respectively. Note that $\mathcal{L}_{t}\cup\mathcal{U}_{t}=[n]$. Then the point that minimizes the uncertainty the most, after it is added to history $H_{t}$, is
$I_{t}=\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{*}\,.$ (1)
This algorithm is a greedy solution to the G-optimal optimal design (Pukelsheim, 2006; Katz-Samuels et al., 2021). Hence we call it G-Optimal design and abbreviated as GO. The pseudocode of GO is in Algorithm 1. We use $u\oplus v$ to denote the concatenation of vectors $u$ and $v$. Note that GO does not depend on observed labels $Y_{t}$. Similar optimal designs have been used to select informative samples in active learning (Chaudhuri et al., 2015; Mukherjee et al., 2022), bandits (Fontaine et al., 2021; Mason et al., 2021; Mukherjee et al., 2023a, b, 2024), and reinforcement learning (Wagenmaker et al., 2022).
The time complexity of GO is $O(d^{2}nT)$. This is because the algorithm searches for the best point out of at most $n$ points in $T$ rounds. The evaluation of each point in each round, $\mathbf{x}_{*}^{\top}(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{*}$, takes $O(d^{2})$ time, because $(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}$ can be implemented in $O(d^{2})$ time using the Sherman-Morrison formula. Note that the LLM is queried only once, in the last step to return $Y_{*}$.
3.2 Simulation-Based Algorithm
Algorithm 2 Simulation-based active learning (SAL)
1:  Input: Target point $\mathbf{x}_{*}$, budget $T$
2:
3:  $\mathcal{L}_{1}\leftarrow\emptyset,\ \mathcal{U}_{1}\leftarrow[n],\ H_{1}\leftarrow()$
4:  for $t=1,\dots,T$ do
5:     for all $i\in\mathcal{U}_{t}$ do
6:        for all $j\in[m]$ do
7:           Sample $Y_{t,i,j}\sim p(\cdot\mid\mathbf{x}_{i},H_{t})$
8:           $H_{t,i,j}\leftarrow H_{t}\oplus(\mathbf{x}_{i},Y_{t,i,j})$
9:           Sample $\tilde{Y}_{t,i,j,1},\tilde{Y}_{t,i,j,2}\sim p(\cdot\mid\mathbf{x}_{*},H_{t,i,j})$
10:        end for
11:     end for
12:     $I_{t}\leftarrow\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2})^{2}$
13:     $X_{t}\leftarrow\mathbf{x}_{I_{t}}$
14:     Observe label $Y_{t}$ of point $X_{t}$
15:     $\mathcal{L}_{t+1}\leftarrow\mathcal{L}_{t}\cup\{I_{t}\},\ \mathcal{U}_{t+1}\leftarrow\mathcal{U}_{t}\setminus\{I_{t}\}$
16:     $H_{t+1}\leftarrow H_{t}\oplus(X_{t},Y_{t})$
17:  end for
18:
19:  Output: Sample $Y_{*}\sim p(\cdot\mid\mathbf{x}_{*},H_{T+1})$
The shortcoming of GO is that the chosen point in round $t$, $X_{t}$, is independent of the previously observed labels, because (1) does not depend on $(Y_{\ell})_{\ell\in[t-1]}$. This is a property of linear models that is undesirable in non-linear models. This motivates our second algorithm design, which simulates the impact of labeling points on uncertainty. Therefore, we call it Simulation-Based Active Learning and abbreviated it as SAL. The pseudocode of SAL is in Algorithm 2.
The key idea in our approach is to replace the closed-form formula in (1) by a simulation. Fix round $t$, history $H_{t}$, and a candidate point $\mathbf{x}_{i}$. To estimate its impact, we simulate its labels $m$ times, by sampling them as $Y_{t,i,j}\sim p(\cdot\mid\mathbf{x}_{i},H_{t})$ for all $j\in[m]$. Let
$H_{t,i,j}=H_{t}\oplus(\mathbf{x}_{i},Y_{t,i,j})$
be the history $H_{t}$ extended by point $\mathbf{x}_{i}$ and its simulated label $Y_{t,i,j}$. Then, to estimate the reduction in uncertainty of $Y_{*}$, we take two independent samples
$\tilde{Y}_{t,i,j,1},\tilde{Y}_{t,i,j,2}\sim p(\cdot\mid\mathbf{x}_{*},H_{t,i,j})$
for each $j\in[m]$ and score point $\mathbf{x}_{i}$ by
$\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2})^{2}\,.$ (2)
The point with the lowest score is chosen and we denote its index by $I_{t}$.
This approach can be justified as follows. Take the same setting as in Section 3.1. Given label $Y_{t,i,j}$ for point $\mathbf{x}_{i}$, the posterior distribution of $\bm{\theta}_{*}\mid H_{t,i,j}$ is $\mathcal{N}(\widehat{\bm{\theta}}_{t,i,j},\widehat{\mathbf{\Sigma}}_{t,i})$, where
$\widehat{\mathbf{\Sigma}}_{t,i}=(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\sigma^{-2}\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}$
is the simulated posterior covariance of $\bm{\theta}_{*}$ and
$\widehat{\bm{\theta}}_{t,i,j}=\widehat{\mathbf{\Sigma}}_{t,i}\left(\widehat{\mathbf{\Sigma}}_{0}^{-1}\bm{\theta}_{0}+\sigma^{-2}\left(\sum_{\ell=1}^{t-1}X_{\ell}Y_{\ell}+\mathbf{x}_{i}Y_{t,i,j}\right)\right)$
is the corresponding posterior mean. By design, $\tilde{Y}_{t,i,j,1}$ and $\tilde{Y}_{t,i,j,2}$ are independent samples from $\mathcal{N}(\mathbf{x}_{*}^{\top}\bm{\theta}_{*},\sigma^{2})$, where $\bm{\theta}_{*}\sim\mathcal{N}(\widehat{\bm{\theta}}_{t,i,j},\widehat{\mathbf{\Sigma}}_{t,i})$. Therefore,
$\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2}\sim\mathcal{N}(0,2(\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}))\,.$
This distribution does not depend on $Y_{t,i,j}$. It also follows that (2) is an unbiased estimate of $2(\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2})$, proportional to $\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}$ in (1). Thus, in linear models, SAL can be viewed as an inefficient implementation of GO.
The time complexity of SAL is $O(nmT)$. This is because the algorithm searches for the best point out of at most $n$ points in $T$ rounds. The evaluation of each point in each round requires $3m$ LLM queries. Therefore, although the time complexity of SAL seems much smaller than that of GO, SAL is much slower in practice.
3.3 Vector Labels
GO and SAL can be easily extended to vector labels. GO does not depend on labels at all. The only modification in SAL is that (2) is replaced with
$\frac{1}{m}\sum_{j=1}^{m}\|\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2})\|_{2}^{2}\,.$ (3)
This is the sum of the posterior variances of label $Y_{*}$ for point $\mathbf{x}_{i}$, over all dimensions of $Y_{*}$.
3.4 Multiple Target Points
We extend GO and SAL to $K$ target points. As a result, they can be applied to classification and regression tasks, where we may want to infer labels of multiple points. We adopt the following notation. The $k$-th target point is $\mathbf{x}_{*,k}$ and its label is $Y_{*,k}$. GO can be extended as follows. In round $t$, the chosen point minimizes the maximum uncertainty of predictions over all target points, which corresponds to
$I_{t}=\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{*,k}\,.$
SAL is modified as follows. Let $\tilde{Y}_{t,i,j,k,1}$ and $\tilde{Y}_{t,i,j,k,2}$ be simulated labels of the LLM for target point $\mathbf{x}_{*,k}$ in round $t$. Then we replace (3) with
$\max_{k\in[K]}\frac{1}{m}\sum_{j=1}^{m}\|\tilde{Y}_{t,i,j,k,1}-\tilde{Y}_{t,i,j,k,2})\|_{2}^{2}\,.$
Finally, both GO and SAL predict labels $Y_{*,k}$ for all target points $\mathbf{x}_{*,k}$ at the end of round $T$.
4 Analysis
In this section, we analyze GO and SAL. The analysis is under the assumption that the labels are scalar. In this case, our objective is to minimize $\mathrm{var}[Y_{*}\mid\mathbf{x}_{*},H_{T+1}]$. The analysis is organized as follows. First, we prove that our objective is decreasing in history but not supermodular, which precludes a straightforward analysis. Second, we analyze GO using the closed form of the posterior covariance $\widehat{\mathbf{\Sigma}}_{t}$. Finally, we prove the equivalence of GO and SAL, and hence provide guarantees for SAL. All analyses are under the assumption of a linear model with Gaussian noise. All proofs are in Appendix B.
4.1 Objective and Its Properties
By the total variance decomposition for a linear model with observation variance $\sigma^{2}$, we get
$\mathrm{var}[Y_{*}\mid\mathbf{x}_{*},H_{T+1}]$
$\quad=\mathrm{var}[\mathbb{E}[Y_{*}\mid\mathbf{x}_{*},\theta_{*},H_{T+1}]\mid\mathbf{x}_{*},H_{T+1}]+{}$
[…]
$\quad=\sigma^{2}\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t}^{-1}\mathbf{x}_{*}+\sigma^{2}\,.$
Therefore, the variance minimization of $Y_{*}\mid\mathbf{x}_{*},H_{T+1}$ is a combinatorial optimization problem,
$\textstyle S_{*}=\mathop{\rm arg\,min}_{S\subseteq[n],|S|=T}f(S)\,,$ (4)
where $f(S)=\mathbf{x}_{*}^{\top}\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{i\in S}\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)^{-1}\mathbf{x}_{*}$ is the optimized function. If $f(S)$ was monotone and supermodular in $S$, we would have $(1-1/e)$-optimality guarantees for GO (Nemhauser et al., 1978). We start with proving the monotonicity of $f(S)$.
Lemma 1.
Function $f(S)$ is monotonically decreasing.
The claim is proved in Section B.1. Now we try to show that $f(S)$ is not supermodular.
Definition 2 (Supermodular function).
Let $[n]$ be a set of $n$ elements. A function $f:2^{[n]}\to\mathbb{R}$ is supermodular if it satisfies the property of diminishing marginal returns. For any $S_{1}\subseteq S_{2}\subseteq[n]$ and $\mathbf{x}\notin S_{2}$, we have $f(S_{1}\cup\left\{\mathbf{x}\right\})-f(S_{1})\leq f(S_{2}\cup\left\{\mathbf{x}\right\})-f(S_{2})$.
Unfortunately $f(S)$ is not supermodular.
Lemma 3.
The function $f(S)$ is not supermodular.
Proof.
Take $\mathbf{\Sigma}_{0}=I_{d}$, $\mathbf{x}_{1}=(1/\sqrt{2},1/\sqrt{2})$, $\mathbf{x}_{2}=(0,1)$, and $\mathbf{x}_{\ast}=(1,0)$. Then $f(\left\{2\right\})-f(\emptyset)=0<0.035714\approx f(\left\{1,2\right\})-f(\left\{1\right\})$. ∎
4.2 Analysis of GO
To overcome Lemma 3, we leverage the properties of the rank-$1$ updates in GO. The proof is under the assumption that the set of points can be partitioned as $\mathcal{X}=S\cup\bar{S}$. The set $S$ represents points that are close to $\mathbf{x}_{*}$. It is convex and defined by $\alpha\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\geq\alpha$ for all $\mathbf{x},\mathbf{y}\in S$. We assume that $S$ is sufficiently large, $|S|\geq T+1$. The set $\bar{S}$ represents points that are not close to $\mathbf{x}_{*}$. It is defined by $\beta\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\leq\beta$ for all $\mathbf{x}\in S$ and $\mathbf{y}\in\bar{S}$. With this in hand, we prove the following claim.
Theorem 4.
Let $\alpha,\beta\geq 0$ be set such that $\beta\geq 1-\alpha^{2}$ and $T\leq\frac{\alpha^{2}}{(\beta+\sqrt{2})\beta d}$. Then
$\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{T+1}\mathbf{x}_{*}\leq\frac{1}{\alpha^{2}T+1}+(1-\alpha^{2})\,.$
The proof is in Section B.2. It is a major departure from similar proofs in active learning with a fixed budget (Azizi et al., 2022; Yang & Tan, 2022) in three aspects. First, we analyze the discrete allocation problem in (4) instead of its continuous optimal-design relaxation (Pukelsheim, 2006). Second, any unlabeled point in $\mathcal{X}$ is labeled at most once. Finally, (4) is asymmetric in the sense that we optimize the uncertainty of a single point $\mathbf{x}_{*}$ over a larger set. To make the analysis manageable, we impose structure on $\mathcal{X}$.
The claim in Theorem 4 holds for any $T$ if $\beta=1/(4dT)$ and $\alpha=\sqrt{1-\beta}$. In this case, $\alpha$ is close to $1$, and we get a near-optimal $O(1/T)$ decrease in posterior variance.
4.3 Analysis of SAL
For a sufficiently large sample size $m$ in SAL, we can establish the following equivalence of SAL and GO.
Theorem 5.
Fix a failure probability $\delta\in(0,1)$. Define $\sigma_{i}^{2}\!\!=\!\!\mathbb{E}[\frac{1}{m}\sum_{j=1}^{m}(\widetilde{Y}_{t,i,j,1}-\widetilde{Y}_{t,i,j,2})^{2}]=2\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}$. Then for any $t\in[T]$ and $i\in[n]$, we have that
$\sigma_{i}^{2}$ $\left[1-2\sqrt{\frac{\log(1/\delta)}{m}}\right]\leq\frac{1}{m}\sum_{j=1}^{m}(\widetilde{Y}_{t,i,j,1}-\widetilde{Y}_{t,i,j,2})^{2}$
$\leq\sigma_{i}^{2}\left[1+2\sqrt{\frac{\log(1/\delta)}{m}}+\frac{2\log(1/\delta)}{m}\right]\,.$
where $\sigma_{i}^{2}\!\!=\!\!\mathbb{E}[\frac{1}{m}\sum_{j=1}^{m}(\widetilde{Y}_{t,i,j,1}-\widetilde{Y}_{t,i,j,2})^{2}]=2\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}$. Moreover, for $m\geq 8\log(1/\delta)$ we have that
$\frac{1}{2}\sigma^{2}_{i}\leq\frac{1}{m}$ $\sum_{j=1}^{m}(\widetilde{Y}_{t,i,j,1}-\widetilde{Y}_{t,i,j,2})^{2}\leq\frac{5}{2}\sigma^{2}_{i}.$
These claims hold with probability at least $1-\delta$.
The claim is proved in Section B.3. The key idea in the proof is that (2) multiplied by $m/[2(\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2})]$ is a chi-squared random variable with $m$ degrees of freedom. Then we use concentration inequalities of Laurent & Massart (2000) to get a high-probability confidence interval on distance to the mean $m$, which in turn allows us to relate the actual variance to its empirical estimate.
Theorem 5 shows that SAL is equivalent to GO for a sufficiently large sample size $m$. Theorem 4 can be then adapted to SAL as follows. The only change is in condition on $T$, which changes to
$T\leq\alpha^{2}\bigg{/}\left(\beta+\sqrt{2}O\left(\tfrac{1-\sqrt{1/m}}{1+\sqrt{1/m}}\right)\right)\beta d+O(\sqrt{1/m})\,.$
Therefore, SAL attains a near-optimal $O(1/T)$ decrease in posterior variance as $m\to\infty$.
5 Experiments
We evaluate GO and SAL on a variety of prediction tasks. These tasks cover both classification and regression, including natural language features, and help us to evaluate the capabilities of GO and SAL to choose few-shot examples for active transductive learning. We also demonstrate that GO and SAL can be used for general pattern recognition. Detailed descriptions of all datasets are in Section C.1.
5.1 Experimental Setup
We use Falcon-40B (Penedo et al., 2023) as the LLM and design prompts following Dinh et al. (2022); Suzgun et al. (2022). Examples of the prompts are given in Appendix D. Each experiment is averaged over 20 runs. At the beginning of each run, we select $K$ target points as the test set where $K$ depends on the task (Section 5.2).
Each run is a simulation that proceeds as follows. In round $t$, each method selects a point $X_{t}$ and observes its true label $Y_{t}$. The compared methods differ only in how this point is chosen. All past interactions $H_{t}=\left\{(X_{\ell},Y_{\ell})\right\}_{\ell\in[t-1]}$ along with the target point $\mathbf{x}_{*,k}$ are used in the LLM prompt to predict the label of $\mathbf{x}_{*,k}$ in round $t$. We denote it by $\tilde{Y}_{*,k,t}$ and use it to evaluate the performance of the method (see Figure 1).
The performance up to round $t$ is evaluated by the cumulative error $\mathcal{L}(t)=\frac{1}{K}\sum_{\ell=1}^{t}\sum_{k=1}^{K}\mathcal{L}(Y_{*,k},\tilde{Y}_{*,k,\ell})$, where $Y_{*,k}$ is the true label of target point $k$, $\tilde{Y}_{*,k,\ell}$ is its predicted label in round $\ell\in[t]$, and $\mathcal{L}(y_{*},y)$ is a task-specific error function. In classification tasks, $\mathcal{L}(y_{*},y)=\mathbb{I}\{y_{*}=y\}$ and we call $\mathcal{L}(t)$ a cumulative misclassification error. In regression tasks, $\mathcal{L}(y_{*},y)=(y_{*}-y)^{2}$ and we call $\mathcal{L}(t)$ a cumulative MSE. Finally, in pattern recognition tasks, $Y_{*,k}$ and $\tilde{Y}_{*,k,\ell}$ are either vectors or matrices. We call $\mathcal{L}(t)$ a cumulative 0-1 error and let $\mathcal{L}(y_{*},y)=\mathbb{I}\{y_{*}=y\}$.
Figure 2: Results on a subset of UCI and OpenML classification and regression tasks. GO and SAL are the best-performing methods in all tasks. Extensive results can be found in Appendix Tables 1 and 2.
The following methods are compared in our experiments:
(1) GO (ours): This is Algorithm 1 where $\mathcal{X}$ are original feature vectors in the dataset.
(2) GO-Inst (ours): This is Algorithm 1 where the original feature vectors are used in the prompt but $\mathcal{X}$ are their $768$-dimensional Instructor embeddings (Su et al., 2022). We use this for natural language features.
(3) SAL (ours): This is Algorithm 2 where $\mathcal{X}$ are original feature vectors in the dataset. To implement SAL efficiently, we combine it with GO as a preprocessing step. Specifically, in round $t$, GO first chooses $5$ most informative points and then we apply SAL. We use $m=1$ in all experiments. We use these approximations since SAL is computationally expensive to run (Section 3.2). Similarly to GO-Inst, we use Instructor embeddings for natural language features.
(4) Uniform: The point in round $t$ is chosen randomly from the unlabeled set $\mathcal{U}_{t}$. Uniform is a pure exploration algorithm that does not take into account the similarity of target points and variance reduction. We choose it as a baseline because it tends to work well in practice. Therefore, it is used frequently in active learning papers (Zhan et al., 2022).
(5) Greedy-NN: The selected point in round $t$ aligns the most with target point $\mathbf{x}_{*}$, $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}\mathbf{x}_{i}$. This baseline shows that our information gathering rule in (1) goes beyond pure feature similarity. For multiple target points $\mathbf{x}_{*,k}$, we use $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\mathbf{x}_{i}$.
(6) D-opt: This baseline is motivated by the D-optimal design (Pukelsheim, 2006; Fedorov, 2013). At round $t$, D-opt computes a distribution over unlabeled points $\mathcal{U}_{t}$, $\pi_{t}=\mathop{\rm arg\,max}_{\pi}\log\operatorname{det}(\sum_{i\in\mathcal{U}_{t}}\pi(i)\mathbf{x}_{i}\mathbf{x}_{i}^{\top})$. The next point is chosen as $I_{t}\sim\pi_{t}$. We choose this baseline to show that other optimal design approaches are not as efficient as GO. Following Kiefer & Wolfowitz (1960); Lattimore & Szepesvári (2020), we note that D-opt can also be viewed as $\pi_{t}=\mathop{\rm arg\,min}_{\pi}\max_{j\in\mathcal{U}_{t}}\mathbf{x}_{j}^{\top}(\sum_{i\in\mathcal{U}_{t}}\pi(i)\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{j}$. This shows the key difference from (1). While GO minimizes the uncertainty of the target point $\mathbf{x}_{*}$ only, D-opt does it for all points in $\mathcal{U}_{t}$.
Figure 3: Results on ARC and PCFG structured prediction tasks. Either GO and SAL is the best-performing method in all tasks.
Figure 4: Results on $4$ synthetic NLP classification tasks. GO and SAL are the best-performing methods in all tasks.
5.2 Results
All used datasets and experimental setups are described in Appendix C. This section only summarizes the main results.
Standard classification and regression tasks. We use $4$ classification and $2$ regression datasets from UCI and OpenML (Section C.1). For classification tasks, $2$ target points per class are chosen from either a single class ($1$-class setting) or two classes ($2$-class setting). For regression tasks, $4$ random target points are used. Our results are reported in Figure 2 and appendix Tables 1 and 2. We observe that GO and SAL are the best-performing methods in all tasks except for 2-class Iris. Larger improvements over the baselines are observed in $1$-class classification and regression tasks. SAL performs slightly better than GO in $1$-class classification tasks. The opposite is observed in the other tasks.
General pattern recognition. We experiment with $4$ tasks: ARC expansion and contraction, ARC rotation, PCFG Add-Subtract, and PCFG Repeat. Both inputs and outputs in these tasks are vectors or matrices. One such example of ARC and PCFG task is shown in Figure 1 and we describe them in detail in Section C.1. Each dataset comprises examples of two patterns: expansion and contraction, clockwise and counter-clockwise rotation, add and subtract, repeat first and second digits. In each run, we choose one randomly as the target point. Our results are reported in Figure 3. In all tasks, GO and SAL are the best-performing methods. SAL outperforms GO in ARC (Alford, 2021; Mirchandani et al., 2023) tasks while GO leads in PCFG (Hupkes et al., 2020).
Natural language classification tasks. We show that GO and SAL work well on general NLP tasks where no explicit numerical features are available. We create three synthetic datasets based on the following tasks: predicting a genre from a movie name (movie-names), predicting a common theme for a pair of movie names (movie-theme), and predicting an entity’s type from its name (entity-names). Each dataset comprises $5$ classes (e.g., there are 5 genres in movie-names: romance, horror, thriller, sport, and action. Details of the other datasets are in Appendix C). In each run, one target point per class is randomly chosen from either a single class or two classes. The feature vectors in GO and SAL are Instructor embeddings of the original text features. Our results are reported in Figure 4. We observe again that GO and SAL are the best-performing methods, although the improvements in the 2-class setting are less significant.
Overall results.To provide a quantitative comparison, we report the mean and standard error of the cumulative error observed at the end of each trial run for all classification tasks (1-class and 2-class), regression tasks, and general pattern recognition tasks. As shown in Appendix Tables 1 and 2, GO and SAL are the absolute top 1 and 2 performing methods in 16 out of 20 tasks. GO ties with Uniform as the top1 on PCFG-repeat, and for the rest 3 tasks, GO and SAL are close to the best-performing method. GO and SAL are comparable, with GO leads in 11 tasks and SAL leads in 8.
Ablation analysis. We hypothesize that GO and SAL perform well because they both reduce the uncertainty of the target point based on the right notion of similarity. To show this, we compare to baselines that only aim to reduce uncertainty uniformly (like Uniform) or only rely on diversity (like D-opt) or only select points with similar features to the target point (Greedy-NN). As shown in our extensive experiments, these baselines fail to match the capabilities of GO and SAL to select informative points. The relative performance of the baselines depends on the task. More extensive ablation studies on Greedy-NN methods comparison can be found in Appendix E. Additionally, we discover that Instructor embedding is mostly effective only on natural language tasks even though it can be used to embed numerical features. As shown in Appendix Figure 8, the accumulative error is worse when Instructor embedding is used on tasks whose original input features are numerical. This is because the similarities are better reflected in the original space, which is important for optimal design.
6 Conclusions
In this paper we studied the framework of active transductive inference (ATI) that systematically chooses the most informative few-shot examples to a given target query. These examples are unlabeled and our framework queries a user to label these examples. These informative labeled examples are then used to minimize the prediction error of the LLM for the given target query. Motivated by the linear model, we proposed an algorithm GO that chooses the most informative examples that minimize the prediction error by minimizing the variance of the posterior co-variance given the target query. We also proposed SAL that uses simulation to estimate the impact of unlabeled examples in reducing the uncertainty of the LLM given a target query. Then it chooses those examples to be labeled that maximally reduce the uncertainty of the LLM for the target query. We analyze these algorithms theoretically and show that they are equivalent in linear models and running these algorithms increases information gain at every iteration. Finally, we show empirically that both GO and SAL result in better prediction accuracy than other baselines in classification, regression, ARC, PCFG, and natural language generation tasks when used with LLMs like Falcon-40B. Our research opens up exciting new directions where we can apply the ATI framework to settings beyond text, such as images and videos using multi-modal LLMs (Yin et al., 2023), or diffusion models (Ho et al., 2020).
Appendix A Related Work
We study the problem of choosing human demonstrations adaptively to get the desired output from the LLM as quickly as statistically possible. We use active learning to choose them and then ask a human to label them. Finally, the human demonstrations are fed as in-context input to the LLM together with the main user query, to obtain the desired output. Thus the name is active transductive inference. Prior works on prompt-tuning and transductive inference (Lester et al., 2021; Dong et al., 2022; Zhang et al., 2022a; Min et al., 2022; Wu et al., 2022; Yu et al., 2022; Suzgun et al., 2022; Liu et al., 2023a; Yu et al., 2023; Liu et al., 2023b) focus on hard prompt-tuning where the user must carefully handcraft the prompt to get the desired output for tasks like movie recommendation, disambiguation QA, navigation, etc. Such examples of carefully handcrafting hard prompts with demonstrations can be found in Suzgun et al. (2022); Srivastava et al. (2022). These papers also show how failing to design such prompts with demonstrations can lead the LLM to predict wrong outputs. Note that we adaptively design the prompt through carefully chosen demonstrations. In our experiments, we show significant improvement over randomly chosen demonstrations.
The problem of active learning is also related to dataset augmentation (Dukler et al., 2021). In this work, the most informative training examples are chosen by optimizing the error of the model on the validation set. The gradient of the validation set error with respect to the weights of training examples has a closed form when the original model is linearized. The main difference in active learning, including in our work, is that labeling all training examples would be costly. Therefore, the labels are not available in advance and are queried adaptively. We also learn in context and do not assume that the gradient information of the LLM is available. Prompt composition has also been an active area of research. Bowman et al. (2023) proposed a-la-carte prompt tuning, where prompts are tuned on individual datasets and composed at inference time to mimic the performance of the model that would have been trained on the union of the corresponding datasets. This idea has been further extended by Perera et al. (2023), where prompts for previously unseen tasks are obtained by linearly combining prompts from known tasks. To do this, they use spectral decomposition and project prompts from known tasks to a lower dimensional space. In our work, we do not tune prompts or compose simpler models. We actively probe an LLM, treated as a black box without any extra side information, to answer a target question as accurately as possible, with as little variance in the answer as possible.
Recently there has been a lot of progress in prompt tuning (or aligning). Hassan et al. (2023) studies prompt aligning for a single test sample to adapt to multi-modal target prompts at test time by minimizing the feature distribution shift to the test domain. In contrast in this paper, we study adapting prompts for many test samples without the feature distribution shift assumption. The (Wang et al., 2023a) trains a smaller LLMto select demonstrations for a larger LLM. However, we rely on active learning to select the smallest number of informative prompts to be labeled by human labelers. This avoids finetuning a smaller LLM for individual tasks. Wang et al. (2023c) studies transductive inference for diffusion models for a different setting where given a pair of task-specific example images, such as depth from/to image and scribble from/to image, and text guidance, the model automatically understands the underlying task and performs the same task on a new query image following the text guidance. However, in our setting, we do not explicitly encode any guidance text. The Wen et al. (2023) proposes to mix the nearest neighbor method with gradient optimization to select prompts during test time. Similarly, Zhang et al. (2023) proposes a nearest neighbor approach to select transductive inference examples for computer vision tasks. We compare our approach against such nearest-neighbor selection algorithms. Finally, Bai et al. (2023); Wang et al. (2023b) analyze transductive inference theoretically to understand its universality, generalization capability, and limitations. In contrast, we only do a theoretical analysis of ATI to show it maximally reduces the estimated variance of the answer to the user’s query. The Zhang et al. (2022b) study the chain-of-thought prompting where they automatically select the prompt using a clustering-based approach.
There are some related works in the area of medical diagnosis chatbot examples that we shared in the introduction. One notable study Caruccio et al. (2024), although not utilizing machine learning techniques, provides valuable insights with its implementation of three hardcoded prompt designs for accurate diagnosis. These designs, however, do not engage in active learning, as they lack the capability to adapt based on user input. Similarly, Kuroiwa et al. (2023) gives more insights into self-diagnostics of orthopedic diseases using ChatGPT. In contrast, web applications such as Buoy Health (BuoyHealth, ) and Live Healthily (livehealthily, ) employ a more dynamic approach, actively tailoring subsequent questions based on users’ responses. This aligns with active learning principles but it is not clear what techniques they apply and is notably underexplored in academic literature, indicating a potential area for further research. Our setting also goes beyond the single shot active prompt tuning studied in Margatina et al. (2023). Note that this work studies prompt tuning only for one iteration, and does not take into account the historical context. So it has limited ability for complex tasks like ARC (Alford, 2021) and PCFG (Hupkes et al., 2020) as well as handling vector labels like GO and SAL.
Active Learning (AL): Recently, there has been a lot of focus on using deep AL to finetune LLMs. All AL algorithms tend to balance uncertainty and diversity in the selection of sample points. We briefly discuss them below and also highlight the main difference of these approaches with prompt aligning with GO and SAL.
(1) Coreset: This is a pure diversity-based approach using a coreset selection. In every iteration, first, the embedding of each point is computed from the network’s penultimate layer, and then sample points are selected using a greedy furthest-first traversal conditioned on all labeled examples (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Citovsky et al., 2021). Observe that in our setting we do not have access to the penultimate layer of the LLM.
(2) Least: This is an uncertainty-based active learning algorithm. Here, the uncertainty score of an example is its predicted class probability. At every iteration, this algorithm then samples points with the smallest uncertainty scores (Settles, 2009, 2011; Wang & Shang, 2014). In our setting, this can be thought of as the sample points closest to the target points. Observe that in our setting Greedy-NNis closest to this baseline, and in Section 5 we show that Greedy-NNis significantly outperformed by GOand SALin the prediction, pcfg, or arc tasks.
(3) Margin: This is also an uncertainty-based active learning algorithm (Tong & Koller, 2001; Balcan et al., 2009; Settles, 2009). At every iteration $t$ it selects samples that are sorted according to their multiclass margin score and then selects points that are the hardest to discriminate and can be thought of as examples closest to their class margin. However, in our setting, we do not have any information on the hypothesis space of the LLM and hence cannot implement such a baseline.
(4) Entropy: This is also an uncertainty-based active learning algorithm (Wang & Shang, 2014; Kremer et al., 2014; Diao et al., 2023). At every iteration $t$ it selects sample points according to the entropy of the example’s predictive class probability distribution. In our setting the Greedy-NN-max-mean is closest to this baseline, and in Figure 7 we show that Greedy-NN-max-mean is outperformed significantly by GO and SAL in the prediction, pcfg, or arc tasks.
(5) Badge: This is an algorithm that combines both uncertainty and diversity sampling (Ash et al., 2019, 2021). For each example $\mathbf{x}$ its gradient embedding $g_{\mathbf{x}}$ is computed with respect to the parameters of the model’s penultimate layer. Finally, Badge chooses a batch of samples to sample by applying $k$-Means++ (Arthur & Vassilvitskii, 2006) on the gradient embeddings. Again recall that we cannot implement such a baseline as we do not have access to the LLMs last layer.
(6) Badge-KM: This algorithm is similar to Badge but in the final step instead of $k$-Means++ it uses $k$-Means on the gradient embeddings. In Yuan et al. (2020) it is observed that applying $k$-Means on the embeddings results in an increase in accuracy over baselines in some datasets. Further Yuan et al. (2020) observed from the t-SNE plots that $k$-Means select centers that are further apart compared to the ones chosen by $k$-Means++ which leads to more diverse sampling in batches.
(7) Bald: Bayesian Active Learning by Disagreements (Kirsch et al., 2019; Gal et al., 2017) chooses examples that are expected to maximize the information gained from the model parameters $\bm{\theta}_{t}$, i.e. the mutual information between predictions and model posterior.
A more comprehensive survey on how AL is used for finetuning deep models can be found in Ren et al. (2021); Zhan et al. (2022). The Bhatt et al. (2024) study how experimental design can be used to select prompts for finetuning a pre-trained LLM. Some recent works have also focused on selecting examples only within a task Wei et al. (2021); Chen et al. (2023); Fifty et al. (2021). However, these works are geared towards selecting prompts within a task for finetuning, whereas we focus on adaptive prompt design using experimental design. The work of Perlitz et al. (2023) also uses AL for finetuning prompts for LLMsto improve label efficiency. The Kung et al. (2023) proposes an AL framework for instruction tuning. However, their approach again focuses on selecting points inside each task and discriminating one task from another. However, they make the simplifying assumption that all points inside the tasks are equally informative which may inhibit the quality of the selected subset.
Appendix B Proofs
This section contains proofs of our main claims.
B.1 Proof of Lemma 1
Recall that the objective function is $f(S)=\mathbf{x}_{*}^{\top}\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{i\in S}\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)^{-1}\mathbf{x}_{*}$. Without loss of generality, let $\sigma^{2}=1$. Fix any $j\in[n]\setminus S$ and let $S_{+}=S\cup\left\{j\right\}$. Then the objective value at $S_{+}$ can be written as
$f(S_{+})=\mathbf{x}_{*}^{\top}\left(\mathbf{\Sigma}_{0}^{-1}+\sum_{i\in S_{+}}\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)^{-1}\mathbf{x}_{*}=\mathbf{x}_{*}^{\top}\left(\mathbf{A}+\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\right)^{-1}\mathbf{x}_{*}\,,$
where $\mathbf{A}=\mathbf{\Sigma}_{0}^{-1}+\sum_{i\in S}\mathbf{x}_{i}\mathbf{x}_{i}^{\top}$. By the Sherman–Morrison formula, we have
$\left(\mathbf{A}+\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\right)^{-1}=\mathbf{A}^{-1}-\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}\,.$
Now note that $\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}$ is a positive semi-definite matrix for any $\mathbf{x}_{j}$ and any positive semi-definite matrix $\mathbf{A}$. As a result, $\mathbf{x}_{*}^{\top}\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}\mathbf{x}_{*}\geq 0$ holds for any $\mathbf{x}_{*}$ and thus
$f(S_{+})=\mathbf{x}_{*}^{\top}\mathbf{A}^{-1}\mathbf{x}_{*}^{\top}-\mathbf{x}_{*}^{\top}\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}\mathbf{x}_{*}\leq\mathbf{x}_{*}^{\top}\mathbf{A}^{-1}\mathbf{x}_{*}^{\top}=f(S)\,.$
This proves our claim.
B.2 Proof of Theorem 4
The proof is under the assumption that points can be partitioned as $\mathcal{X}=S\cup\bar{S}$. The set $S$ represents points that are close to $\mathbf{x}_{*}$. It is convex and defined by $\alpha\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\geq\alpha$ for all $\mathbf{x},\mathbf{y}\in S$. The set $\bar{S}$ represents points that are not close to $\mathbf{x}_{*}$. It is defined $\beta\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\leq\beta$ for all $\mathbf{x}\in S$ and $\mathbf{y}\in\bar{S}$. The inverse of the covariance matrix after $t-1$ observations is
$\widehat{\mathbf{\Sigma}}_{t}^{-1}=\mathbf{\Lambda}_{t}=\mathbf{I}_{d}+\sum_{\ell=1}^{t-1}\mathbf{x}_{\ell}\mathbf{x}_{\ell}^{\top}=\sum_{i=1}^{d}\lambda_{t,i}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\,.$
The latter is the eigendecompositon of $\mathbf{\Lambda}_{t}$, where $\lambda_{t,i}$ is the $i$-th largest eigenvalue and $\mathbf{v}_{t,i}$ is the corresponding eigenvector. Note that $\mathbf{\Lambda}_{t}^{-1}=\sum_{i=1}^{d}\lambda_{t,i}^{-1}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}$. To simplify exposition, we assume that all points have unit length. We analyze the eigenvalues of $\mathbf{\Lambda}_{t}$ first.
Lemma 6.
For all $i\in[d]$, $1\leq\lambda_{t,i}\leq t$. Moreover, let $\mathbf{x}_{\ell}\in S$ hold for all $\ell\in[t-1]$. Then $\lambda_{t,1}\geq\alpha^{2}(t-1)+1$.
Proof.
The first claims follows directly from the definition of $\mathbf{\Lambda}_{t}$ and that $\|\mathbf{x}_{\ell}\|_{2}\leq 1$. The second claim is proved using the definition of the maximum eigenvalue,
$\lambda_{t,1}=\mathbf{v}_{t,1}^{\top}\mathbf{\Lambda}_{t}\mathbf{v}_{t,1}\geq\mathbf{x}_{*}^{\top}\left(\mathbf{I}_{d}+\sum_{\ell=1}^{t-1}\mathbf{x}_{\ell}\mathbf{x}_{\ell}^{\top}\right)\mathbf{x}_{*}\geq\alpha^{2}(t-1)+1\,.$
The last inequality follows from $\mathbf{x}_{*}^{\top}\mathbf{x}_{\ell}\geq\alpha$. This completes the proof. ∎
We continue with claims about the eigenvectors of $\mathbf{\Lambda}_{t}$.
Lemma 7.
Let $\mathbf{x}_{\ell}\in S$ hold for all $\ell\in[t-1]$. Then $\mathbf{v}_{t,1}\in S$. Moreover, let $\beta\geq 1-\alpha^{2}$. Then $\mathbf{v}_{t,i}\in\bar{S}$ for all $i\geq 2$.
Proof.
Since all $\mathbf{x}_{\ell}\in S$ and $S$ is a convex set, $\mathbf{v}_{t,1}\in S$. Now take any $\mathbf{x}\in S$ and $i\geq 2$, and note that
$(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}\leq\sum_{i=2}^{d}(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}=1-(\mathbf{x}^{\top}\mathbf{v}_{t,1})^{2}\leq 1-\alpha^{2}\,.$
We use that $\sum_{i=1}^{d}(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}=1$ and $\mathbf{x}^{\top}\mathbf{v}_{t,1}\geq\alpha$. Therefore, when $\beta\geq 1-\alpha^{2}$, we have $\mathbf{v}_{t,i}\in\bar{S}$ for all $i\geq 2$. ∎
Our analysis has two parts. First, we bound the approximation error under the assumption that $\mathbf{x}_{\ell}\in S$ holds for all $\ell\in[T]$. Second, we show how to choose $\alpha$ and $\beta$ to guarantee this. We start with the approximation error.
Lemma 8.
Let $\mathbf{x}_{\ell}\in S$ hold for all $\ell\in[T]$. Then
$\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{T+1}^{-1}\mathbf{x}_{*}\leq\frac{1}{\alpha^{2}T+1}+(1-\alpha^{2})\,.$
Proof.
We start with
$\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{T+1}^{-1}\mathbf{x}_{*}=\sum_{i=1}^{d}\lambda_{T+1,i}^{-1}\mathbf{x}_{*}^{\top}\mathbf{v}_{T+1,i}\mathbf{v}_{T+1,i}^{\top}\mathbf{x}_{*}\leq\frac{(\mathbf{x}_{*}^{\top}\mathbf{v}_{T+1,1})^{2}}{\alpha^{2}T+1}+\sum_{i=2}^{d}(\mathbf{x}_{*}^{\top}\mathbf{v}_{T+1,i})^{2}\,.$
The inequality uses lower bounds in Lemma 6. Then we apply $\mathbf{x}_{*}^{\top}\mathbf{v}_{T+1,1}\leq 1$ and $\sum_{i=2}^{d}(\mathbf{x}_{*}^{\top}\mathbf{v}_{T+1,i})^{2}\leq 1-\alpha^{2}$. ∎
Now we prove by induction that $\mathbf{x}_{\ell}\in S$ holds for all $\ell\in[T]$.
Lemma 9.
Let $\mathbf{x}_{\ell}\in S$ hold for all $\ell\in[t-1]$. Suppose that $t\leq\frac{\alpha^{2}}{(\beta+\sqrt{2})\beta d}$. Then $\mathbf{x}_{t}\in S$.
Proof.
Our algorithm chooses a point $x_{t}\in S$ when
$\frac{\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*}}{1+\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}}\geq\frac{\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*}}{1+\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}}$
holds for any $\mathbf{x}\in S$ and $\mathbf{y}\in\bar{S}$. Since $0\leq\mathbf{v}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{v}\leq 1$ for any $\|\mathbf{v}\|_{2}\leq 1$, the above event occurs when
$(\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x})^{2}=\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*}\geq 2\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*}=2(\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y})^{2}\,.$
We start with an upper bound on the right-hand side,
$|\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}|\leq\sum_{i=1}^{d}\lambda_{t,i}^{-1}|\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\mathbf{y}|\leq\beta d\,.$
Here we use $\lambda_{t,i}\geq 1$ (Lemma 6), and that $\mathbf{v}_{t,1}^{\top}\mathbf{y}\leq\beta$ and $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\leq\beta$ when $i\geq 2$.
Now we bound the left-hand side as
$|\mathbf{x}_{*}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}|\geq\lambda_{t,1}^{-1}|\mathbf{x}_{*}^{\top}\mathbf{v}_{t,1}\mathbf{v}_{t,1}^{\top}\mathbf{x}|-\sum_{i=2}^{d}\lambda_{t,i}^{-1}|\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\mathbf{x}|\geq\frac{\alpha^{2}}{t}-\beta^{2}d\,.$
To bound the first term, we use $\lambda_{t,1}\leq t$, and that $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,1}\geq\alpha$ and $\mathbf{v}_{t,1}^{\top}\mathbf{x}\geq\alpha$. To bound the second term, we use $\lambda_{t,i}\geq 1$, and that $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\leq\beta$ and $\mathbf{v}_{t,i}^{\top}\mathbf{x}\leq\beta$.
Now we chain all inequalities and get that our algorithm chooses a point $x_{t}\in S$ when
$t\leq\frac{\alpha^{2}}{(\beta+\sqrt{2})\beta d}\,.$
This completes the proof. ∎
B.3 Proof of Theorem 5
Consider a single test point $\mathbf{x}_{*}$. Fix $m=1$. Recall that $H_{t}=\left(X_{\ell},Y_{\ell}\right)_{\ell\in[t-1]}$ is the history till round $t$. Then the posterior variance is
$\widehat{\mathbf{\Sigma}}_{t}=\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}\right)^{-1}$
and $\widehat{\bm{\theta}}_{t}=\widehat{\mathbf{\Sigma}}_{t}\left(\mathbf{\Sigma}_{0}^{-1}\bm{\theta}_{0}+\sigma^{-2}\sum_{\ell=1}^{t-1}X_{\ell}Y_{\ell}\right)$. Now fix a $X_{t}$ at round $t$. Then $Y_{t}=X_{t}^{\top}\bm{\theta}_{*}+\epsilon_{t}$ where $\bm{\theta}_{*}\mid H_{t}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t},\widehat{\mathbf{\Sigma}}_{t}\right)$ and $\epsilon_{t}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. Then $\bm{\theta}_{*}|H_{t+1}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t+1},\widehat{\mathbf{\Sigma}}_{t+1}\right)$ holds for any $Y_{t}$.
Now fix an $\mathbf{x}_{i}\in\mathcal{X}$ and add it to $\widehat{\mathbf{\Sigma}}_{t}$. Define $\hat{\sigma}_{t,i}^{2}=\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2})^{2}$ and the $\mathbb{E}[\hat{\sigma}_{t,i}^{2}]=\sigma_{i}^{2}$. Then using Lemma 10 we can show that with probability $(1-\delta)$
$\sigma_{i}^{2}\left[1-2\sqrt{\frac{\log(1/\delta)}{m}}\right]\leq\hat{\sigma}_{t,i}^{2}\leq\sigma_{i}^{2}\left[1+2\sqrt{\frac{\log(1/\delta)}{m}}+\frac{2\log(1/\delta)}{m}\right]$ (5)
Set $m\geq 8\log(1/\delta)$ in (5). It follows then that
$\frac{1}{2}\sigma^{2}_{i}\leq\frac{1}{m}$ $\sum_{j=1}^{m}(\widetilde{Y}_{t,i,j,1}-\widetilde{Y}_{t,i,j,2})^{2}\leq\frac{5}{2}\sigma^{2}_{i}.$
Hence, to minimize the quantity $\widetilde{Y}_{t,i,j,1}-\widetilde{Y}_{t,i,j,2}$ we should be minimizing the variance $2\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}$. Observe that minimizing the variance in SAL leads to minimizing the quantity $2\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}$ which is same as minimizing the score $\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}$ for GO. The claim of the theorem follows.
Lemma 10.
Fix round $t\in[T]$, a sample point $\mathbf{x}_{i}$, and a test point $\mathbf{x}_{*}$, and failure probability $\delta\in(0,1)$. Let
$m=m_{t,i}-1$
and suppose that $m>4\log(1/\delta)$. Define $\hat{\sigma}_{t,i}^{2}=\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}_{t,i,j,1}-\tilde{Y}_{t,i,j,2})^{2}$ and the $\mathbb{E}[\hat{\sigma}_{t,i}^{2}]=\sigma_{i}^{2}$. Then
$\mathbb{P}\left(\hat{\sigma}_{t,i}^{2}\leq\sigma_{i}^{2}\left[1-2\sqrt{\frac{\log(1/\delta)}{m}}\right]\right)\leq\delta$
holds with probability at least $1-\delta$. Analogously,
$\mathbb{P}\left(\hat{\sigma}_{t,i}^{2}\geq\sigma_{i}^{2}\left[1+2\sqrt{\frac{\log(1/\delta)}{m}}+\frac{2\log(1/\delta)}{m}\right]\right)\leq\delta$
holds with probability at least $1-\delta$.
Proof.
Fix an $\mathbf{x}_{i}\in\mathcal{X}$ and add it to $\widehat{\mathbf{\Sigma}}_{t}$. Denote this new co-variance matrix as $\widehat{\mathbf{\Sigma}}_{t,i}$. Let $\widetilde{Y}_{t,i,j,1}=\mathbf{x}_{*}^{\top}\bm{\theta}_{*}+\epsilon_{t,i,j,1}$, where $\bm{\theta}_{*}\mid H_{t}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t,i},\widehat{\mathbf{\Sigma}}_{t+1}\right)$ and $\epsilon_{t,i,j,1}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. This yields that $\widetilde{Y}_{t,i,j,1}\sim\mathcal{N}\left(\mathbf{x}_{*}^{\top}\widehat{\bm{\theta}}_{t+1},\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}\right)$.
Similarly $\widetilde{Y}_{t,i,j,2}=\mathbf{x}_{*}^{\top}\bm{\theta}_{*}+\epsilon_{t,i,j,2}$, where $\bm{\theta}_{*}\mid H_{t}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t+1},\widehat{\mathbf{\Sigma}}_{t,i}\right)$ and $\epsilon_{t,i,j,2}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. Therefore, we get that
$\widetilde{Y}_{t,i,j,1}-\widetilde{Y}_{t,i,j,2}\sim\mathcal{N}(0,2\mathbf{x}_{*}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*}+\sigma^{2}).$
Now we proceed the same way as in Lemma 2 of Lalitha et al. (2023). Using Cochran’s theorem, we have that $\hat{\sigma}_{t,i}^{2}m/\sigma_{i}^{2}$ is a $\chi^{2}$ random variable with $m$ degrees of freedom. Then using (4.4) and Lemma 1 of Laurent & Massart (2000) we can show that
$\mathbb{P}\left(m-\frac{\hat{\sigma}_{s,t,i}^{2}m}{\sigma_{i}^{2}}\geq 2\sqrt{m\log(1/\delta)}\right)\leq\delta$ (6)
Dividing both sides of (6) in the probability by $m$, and multiplying by $\sigma_{i}^{2}$, we can get the following
$\mathbb{P}\left(\sigma_{i}^{2}(1-2\sqrt{\log(1/\delta)/m})\geq\hat{\sigma}_{s,t,i}^{2}\right)\leq\delta.$
Observe that $01-2\sqrt{\log(1/\delta)/m}>$ 0, we can divide both sides by it and get the first claim of the lemma. The second claim is proved by (4.3) in Laurent & Massart (2000), an immediate corollary of their Lemma 1, we have
$\mathbb{P}\left(\frac{\hat{\sigma}_{s,t,i}^{2}m}{\sigma_{i}^{2}}-m\geq 2\sqrt{m\log(1/\delta)}+2\log(1/\delta)\right)\leq\delta$
The claim of the lemma follows. ∎
Appendix C Additional Experiments and Results
We now briefly discuss the various datasets used in this work.
C.1 Datasets
We now briefly describe the datasets used for our experiments. All the real-life datasets are from UCI (Markelle Kelly, 1988) and OpenML (Vanschoren et al., 2013) repositories. We use $4$ classification and $3$ regression datasets from UCI and OpenML. Additionally, we use $2$ custom datasets for movie names and entity names for classification task in our experiments. These are as follows:
(1) Iris: We use this UCI dataset for classification task. This dataset consists of four features of flowers and three classes of flowers. We use all four features in the prompts as well as estimating the score for selecting the next action.
(2) Banknote-authentication: We use this OpenML dataset for classification task. This dataset consists of five features of banknotes and two classes for identifying fake or original banknote. Out of these five features, we use four features in the prompts as well as estimating the score for selecting the next action.
(3) Balance-scale: We use this OpenML dataset for classification task. This dataset consists of five features of a scale and three classes of whether the scale tips left/right or is balanced. Out of these five features, we use four features in the prompts as well as estimating the score for selecting the next action.
(4) Thyroid-new: We use this OpenML dataset for classification task. This dataset consists of six features for thyroids and three classes. Out of these six features, we use five features in the prompts as well as estimating the score for selecting the next action.
(5) Movie-name: We use this custom dataset for classification task. This dataset consists of movie names across five genres (classes) romance, horror, thriller, sport, and action. We convert the movie names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the movie names and the goal is to identify the common genre.
(6) Movie-theme: We use this custom dataset for classification tasks in identifying a common theme between pairs of movies. This dataset consists of movie names across five themes (classes) good-vs-evil, man-vs-nature, redemption, Love conquers all, and coming-of-age. We convert the movie names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the pair of movie names and the goal is to identify the common theme.
(7) Entity-name: We use this custom dataset for classification task. This dataset consists of entity names across five entity types (classes) like mountains, seas, rivers, vehicles, and celebrities. Again, we convert the entity names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the entity names and the goal is to identify the entity type.
(8) Fifa: We use this OpenML dataset for the regression task. This dataset consists of six features of players and the clubs they joined as targets. Out of these six features, we use five features in the prompts as well as estimating the score for selecting the next action.
(9) Machine-cpu: We use this OpenML dataset for regression tasks. This dataset consists of seven features of machine cpu and the target variable is the performance of the cpu. Out of these seven features, we use five features in the prompts as well as estimating the score for selecting the next action.
C.1.1 ARC Experiment
In the recent works of Mirchandani et al. (2023); Srivastava et al. (2022) they showed that LLMs behave as general pattern matching machine. In fact they showed that LLMs can be used to solve tasks from Abstract Reasoning Corpus (ARC) tasks. In the following experiments we choose two such tasks: (1) ARC expansion and contraction experiment and (2) ARC rotation experiment.
(1) ARC expansion and contraction experiment: In the expansion and contraction experiment, there are two sets of input matrices of dimension $4\times 4$ which constitute the input feature space $\mathcal{X}$. The first set of input matrices are have integer values in their center $2\times 2$ cells while all the other cells are 00. The output space $\mathcal{Y}$ of this $4\times 4$ matrix is also a $4\times 4$ matrix where the $4$ inner cells have moved to the $4$ corners. These matrices are termed as expansion matrices.
Similarly, the other set of $4\times 4$ matrices have the $4$ non-zeros values in their corners. These constitute the input feature space $\mathcal{X}$. Then the output space $\mathcal{Y}$ is given by $4\times 4$ matrix where the four non-zeros cells come to the center and all the other cell values are 00. These matrices are termed as contraction matrices. This is shown in Figure 5(a).
Therefore, the input feature space consist of both the expansion and contraction matrices. At every trial the target point is chosen randomly from any one of the target set. Then we run the four baselines for $20$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Figure 5(c).
(2) ARC rotation experiment: In the rotation experiment, there are again two sets of input matrices of dimension $4\times 4$ which constitute the input feature space $\mathcal{X}$. The first set of input matrices are have integer values in their four corner cells while all the other cells are 00. The output space $\mathcal{Y}$ of this $4\times 4$ matrix is also a $4\times 4$ matrix where the $4$ corner cells values have moved $90^{\circ}$ in the clockwise direction. These matrices are termed as clockwise matrices.
Similarly, the other set of $4\times 4$ matrices have the $4$ non-zeros values in their corners. These constitute the input feature space $\mathcal{X}$. Then the output space $\mathcal{Y}$ is given by $4\times 4$ matrix where the four non-zeros cells have moved $15^{\circ}$ in the anti-clockwise direction and all the other cell values are 00. These matrices are termed as anit-clockwise matrices. This is shown in Figure 5(b).
Therefore, the input feature space consist of both the clockwise and anti-clockwise matrices. At every trial the target point is chosen randomly from any one of the target set. Then we run the four baselines for $20$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Figure 5(d).
(a) Classification on ARC Expansion and Contraction Experiment
(b) Classification on ARC Rotation Experiment
(c) Classification on Expansion and Contraction
(d) Classification on ARC Rotation Experiment
Figure 5: Experiment results with Falcon 40B on ARC tasks. For the classification tasks, we show the cumulative mis-classification error in the y-axis.
C.1.2 PCFG Experiment
In this experiment the goal is to predict the next output of a sequence. In the following experiments we choose two such tasks: (1) PCFG add-subtract experiment and (2) PCFG repeat experiment.
(1) PCFG add-subtract experiment: In the add-subtract experiment, there are two sets of input sequences of $4$ integers which constitute the input feature space $\mathcal{X}$. The first set of input sequences consist of odd integer values. The output space $\mathcal{Y}$ of this sequences of $4$ integers is a sequences of $5$ integers where the last integer is padded to the original sequence by adding one to the last odd integer. These sequences are termed as add sequences.
Similarly, the other set of input sequences consist of even integer values. The output space $\mathcal{Y}$ of this sequences of $4$ even integers is a sequences of $5$ integers where the last integer is padded to the original sequence by subtracting one from the last even integer. These sequences are termed as even sequences. This is shown in Figure 6(a).
Therefore, the input feature space consist of both the odd and even sequences. At every trial the target point is chosen randomly from any one of the target set. Then we run the four baselines for $20$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Figure 6(c).
(1) PCFG repeat experiment: In the repeat experiment, there are two sets of input sequences of $4$ integers which constitute the input feature space $\mathcal{X}$. The first set of input sequences consist of odd integer values. The output space $\mathcal{Y}$ of this sequences of $4$ integers is a sequences of $5$ integers where the last integer is padded to the original sequence by repeating the first odd integer. These sequences are termed as odd-repeat sequences.
Similarly, the other set of input sequences consist of even integer values. The output space $\mathcal{Y}$ of this sequences of $4$ even integers is a sequences of $5$ integers where the last integer is padded to the original sequence by repeating the second even integer. These sequences are termed as even-repeat sequences. This is shown in Figure 6(b).
Therefore, the input feature space consist of both the odd-repeat and even-repeat sequences. At every trial the target point is chosen randomly from any one of the target set. Then we run the four baselines for $20$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Figure 6(d).
(a) Classification on PCFG add-subtract Experiment
(b) Classification on PCFG repeat Experiment
(c) Classification on PCFG add-subtract Experiment
(d) Classification on PCFG repeat Experiment
Figure 6: Experiment results with Falcon 40B on PCFG tasks. For the classification tasks, we show the cumulative mis-classification error in the y-axis.
C.2 Additional results
To provide a more quantitative comparison, we compute the mean and standard error of the cumulative error observed at the end of each trial run. As shown in Tables 1 and 2, GO and SAL are the absolute top 1 and 2 performing methods in 16 out of 20 tasks, and are close to the best-performing method on the rest 4 tasks.
Table 1: Cumulative error at the end of each run. The mean and standard errors for 20 repeat runs are shown for all classification datasets for 1-class and 2-class settings, with the top-1 method in bold. For movite-name, movie-theme, and entity-name, Instructor embedding is used as the feature for active experiment design.SAL and GO outperforms other baselines on the majority of the tasks.
Dataset iris banknote balance thyroid movie-name movie-theme entity-name
1-class classification
Uniform $9.05\pm 0.81$ $10.95\pm 0.45$ $12.40\pm 0.54$ $11.53\pm 0.80$ $12.05\pm 0.56$ $15.07\pm 0.57$ $13.40\pm 0.57$
Greedy-NN $13.43\pm 0.35$ $10.57\pm 0.29$ $12.75\pm 0.77$ $12.05\pm 0.53$ $13.18\pm 0.89$ $12.12\pm 1.17$ $15.07\pm 0.70$
D-opt $10.03\pm 0.76$ $10.22\pm 0.38$ $13.15\pm 0.63$ $12.93\pm 0.71$ $11.55\pm 0.54$ $15.53\pm 0.56$ $13.43\pm 0.48$
GO (ours) $7.12\pm 1.01$ $8.32\pm 0.71$ $9.78\pm 0.97$ $\mathbf{7.47\pm 0.80}$ $8.40\pm 0.54$ $\mathbf{11.20\pm 0.82}$ $\mathbf{9.32\pm 0.78}$
SAL (ours) $\mathbf{7.05\pm 0.93}$ $\mathbf{8.25\pm 0.74}$ $\mathbf{9.03\pm 0.91}$ $7.88\pm 0.43$ $\mathbf{8.25\pm 0.56}$ $11.55\pm 0.83$ $10.15\pm 0.76$
2-class classification
Uniform $\mathbf{8.79\pm 0.53}$ $10.57\pm 0.34$ $12.86\pm 0.43$ $12.19\pm 0.53$ $11.88\pm 0.57$ $14.51\pm 0.53$ $12.66\pm 0.40$
Greedy-NN $13.43\pm 0.27$ $10.72\pm 0.23$ $13.75\pm 0.39$ $12.41\pm 0.25$ $13.86\pm 0.63$ $\mathbf{13.80\pm 0.69}$ $14.35\pm 0.44$
D-opt $8.95\pm 0.52$ $10.51\pm 0.26$ $13.16\pm 0.35$ $11.47\pm 0.35$ $11.29\pm 0.44$ $14.55\pm 0.56$ $\mathbf{11.54\pm 0.46}$
GO (ours) $8.93\pm 0.50$ $10.80\pm 0.28$ $\mathbf{11.79\pm 0.25}$ $\mathbf{10.75\pm 0.45}$ $11.55\pm 0.48$ $14.06\pm 0.39$ $11.69\pm 0.35$
SAL (ours) $8.93\pm 0.49$ $\mathbf{10.40\pm 0.18}$ $12.75\pm 0.33$ $11.12\pm 0.37$ $\mathbf{11.18\pm 0.68}$ $14.25\pm 0.36$ $12.29\pm 0.49$
Table 2: Cumulative error at the end of each run. Mean and standard error for 20 repeat runs are shown for all ARC, PCFG and regression datasets, with top-1 method in bold. SAL and GO outperform other baselines on the majority of the tasks.
Dataset arc_expand_contract arc_rotation pcfg_add_subtract pcfg_repeat fifa CPU
Uniform $3.60\pm 0.33$ $4.25\pm 0.19$ $14.25\pm 0.70$ $\mathbf{6.25\pm 0.83}$ $31.14e4\pm 4.36e$ 3 $33.67e4\pm 3.61e$ 3
Greedy-NN $2.55\pm 0.38$ $5.00\pm 0.00$ $19.85\pm 0.08$ $18.65\pm 0.31$ $42.39e4\pm 1.15e$ 4 $41.08e5\pm 1.37e$ 4
D-opt $3.90\pm 0.27$ $4.15\pm 0.21$ $14.45\pm 0.78$ $9.90\pm 0.74$ $32.74e4\pm 3.36e$ 3 $32.99e4\pm 2.62e$ 3
GO (ours) $1.50\pm 0.26$ $2.35\pm 0.30$ $\mathbf{10.70\pm 0.68}$ $\mathbf{6.25\pm 0.72}$ $\mathbf{6.78e3\pm 2.07e3}$ $\mathbf{1.75e4\pm 2.05e3}$
SAL (ours) $\mathbf{1.00\pm 0.21}$ $\mathbf{1.85\pm 0.18}$ $11.35\pm 0.93$ $6.90\pm 0.61$ $38.80e3\pm 2.99e$ 3 $31.77e4\pm 2.16e$ 3
Appendix D Prompt Examples
Iris Classification Dataset: Below we give an example of how we use the prompts to be used in the LLM for the Iris misclassification task. Similar types of prompt can be found in Dinh et al. (2022); Suzgun et al. (2022). In the following prompt, the point $\mathbf{x}_{i}$ chosen by GO at trial $1$, time $1$ is given by $[42,29,69,75]\in\mathbb{R}^{4}$. The label of this point is done by human and is the class of this data point (shown as $2$). Finally, this prompt is fed into the LLM as $x1=42,x2=29,x3=69,x4=75$ for the target query $\mathbf{x}_{*,1}$ represented as $x1=69,x2=33,x3=64,x4=54$ for the LLM prediction of the class of the target query. Finally observe that we proceed in the similar way for each of the $\mathbf{x}_{*,2}$ as $x1=19,x2=58,x3=8,x4=4$, $\mathbf{x}_{*,2}$ as $x1=19,x2=62,x3=5,x4=8$, and $\mathbf{x}_{*,4}$ as $x1=50,x2=25,x3=78,x4=54$. Then we move to the next iteration of the trail, and this interaction proceeds till the end of the horizon.
trial 1
action in time 1: [42 29 69 75]
final_prompt at time 1:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 69, x2 = 33, x3 = 64, x4 = 54 label is
answer in time 1: 0
True label: 1
final_prompt at time 1:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 19, x2 = 58, x3 = 8, x4 = 4 label is
answer in time 1: 1
True label: 0
final_prompt at time 1:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 19, x2 = 62, x3 = 5, x4 = 8 label is
answer in time 1: 0
True label: 0
final_prompt at time 1:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 42, x2 = 33, x3 = 69, x4 = 96 label is
answer in time 1: 1
True label: 2
final_prompt at time 1:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 69, x2 = 33, x3 = 64, x4 = 54 label is
answer in time 1: 1
True label: 1
final_prompt at time 2:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 50, x2 = 25, x3 = 78, x4 = 54 label is
answer in time 2: 1
True label: 2
action in time 2: [56 12 58 50]
final_prompt at time 2:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 56, x2 = 12, x3 = 58, x4 = 50 label is 1. x1 = 69, x2 = 33, x3 = 64, x4 = 54 label is
answer in time 2: 0
True label: 1
final_prompt at time 2:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 56, x2 = 12, x3 = 58, x4 = 50 label is 1. x1 = 19, x2 = 58, x3 = 8, x4 = 4 label is
answer in time 2: 3
True label: 0
final_prompt at time 2:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 56, x2 = 12, x3 = 58, x4 = 50 label is 1. x1 = 19, x2 = 62, x3 = 5, x4 = 8 label is
answer in time 2: 0
True label: 0
final_prompt at time 2:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 56, x2 = 12, x3 = 58, x4 = 50 label is 1. x1 = 42, x2 = 33, x3 = 69, x4 = 96 label is
answer in time 2: 1
True label: 2
[…]
trial 1
action in time 1: [’Frozen (2013)’ ’Brave (2012)’]
final_prompt at time 1:
[’Frozen (2013)’ ’Brave (2012)’] label is 2, Monsters Inc, Frozen (2013):
answer in time 1: 2
True label: 2
action in time 2: [’Wall’ ’Monsters Inc’]
final_prompt at time 2:
[’Wall’ ’Monsters Inc’] label is 2, Monsters Inc, Frozen (2013):
answer in time 2: 2
True label: 2
[…]
trial 1
action in time 1: [’Monty Python and the Holy Grail (1975)’]
final_prompt at time 1:
Monty Python and the Holy Grail (1975) label is 4, Moneyball (2011):
answer in time 0: 4
 True label: 1
final_prompt at time 1:
Monty Python and the Holy Grail (1975) label is 4, Slap Shot (1977):
answer in time 1: 4
 True label: 1
Appendix E Ablation Studies
E.1 Extensive ablation experiments with Greedy-NN
In this section we conduct an ablation study for various versions of Greedy-NN.
Greedy-NN-max-max: This is the version implemented in the paper in all experiments which we call as Greedy-NN. For clarity in this section, we will call this as Greedy-NN-max-max. Recall the Greedy-NN-max-max implemented in Section 5.1 is as follows: The chosen point in round $t$ aligns the most with the target point $\mathbf{x}_{*}$ such that $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}\mathbf{x}_{i}$. For multiple target points $\mathbf{x}_{*,k}$ Greedy-NN-max-max chooses $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\max_{k}\mathbf{x}_{*,k}^{\top}\mathbf{x}_{i}$.
Greedy-NN-max-min: The Greedy-NN-max-min works as follows: The chosen point in round $t$ aligns the most with the target point $\mathbf{x}_{*}$ such that $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}\mathbf{x}_{i}$. For multiple target points $\mathbf{x}_{*,k}$ Greedy-NN-max-min chooses $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\min_{k}\mathbf{x}_{*,k}^{\top}\mathbf{x}_{i}$.
Greedy-NN-max-mean: Before we state this method, we define the notation $\mathrm{avg}(\mathbf{x}_{i})=\frac{1}{K}\sum_{k=1}^{K}\mathbf{x}_{*,k}^{\top}\mathbf{x}_{i}$. Then define the score for a point $\mathbf{x}_{i}$ with respect to $\mathbf{x}_{*}$ and the distance from the average $\mathrm{avg}(\mathbf{x}_{i})$ as $s(\mathbf{x}_{i},\mathbf{x}_{*})=|\mathbf{x}_{*}^{\top}\mathbf{x}_{i}-\mathrm{avg}(\mathbf{x}_{i})|$. Now, the Greedy-NN-max-mean works as follows: At round $t$ choose a point that aligns most with the target point $\mathbf{x}_{*}$. So the point is chosen such that $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}\mathbf{x}_{i}$. For multiple target points $\mathbf{x}_{*,k}$ Greedy-NN-max-mean chooses $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\mathop{\rm arg\,min}_{k}s(\mathbf{x}_{i},\mathbf{x}_{*,k})$. Observe that the inner $\min$ chooses the point $\mathbf{x}_{i}$ closest to the average of all the target points $\mathbf{x}_{*,k}$ to balance exploration.
(a) Classification on ARC Expansion and Contraction Experiment
(b) Classification on ARC Rotation Experiment
(c) Classification on PCFG Add Subtract
(d) Classification on Iris
Figure 7: Experiment results with Falcon 40B on ARC, PCFG, classification tasks. For the classification tasks, we show the cumulative mis-classification error in the y-axis.
The experiments with all these variants are shown in Figure 7. From this Figure 7 we see that none of the variants of the Greedy-NN are reducing the misclassification error or 0-1 error as fast as GO or SAL. This performance of the many variants of Greedy-NN can be attributed to its lack of correctly balancing exploration-exploitation.
E.2 Instructor embedding vs numerical features
Instructor embedding can be derived for any inputs that can be represented as strings. In this section, we examine whether Instructor embedding is helpful in active experiment design when numerical input features are considered. As shown in Figure 8, using Instructor embedding for the numerical features always hurts the performance of G-optimal active experiment design.
Figure 8: Comparison of the cumulative error for active experiment design conducted with original numerical features vs. with Instructor embedding. The selected datasets all use numerical input features rather than natural language. GO is used in all the experiments in this figure.
Appendix F Table of Notations
Notations Definition
$n$ Total unlabeled points
$d$ Dimension of the feature
$\mathcal{X}$ Feature set
$\mathcal{Y}$ Label space
$\bm{\theta}_{*}$ Unknown model parameter
$\mathbf{x}_{i}$ Feature of sample point $i$
$\mathbf{x}_{*,k}$ $k$-th target point
$f(\mathbf{x},\bm{\theta}_{*})$ Model
$Y_{*}$ Target label
$H_{t}=(X_{\ell},Y_{\ell})_{\ell\in[t-1]}$ History of $t-1$ previously labeled points
$p(\cdot\mid\mathbf{x},H_{t})$ Distribution of the label of point $\mathbf{x}$ conditioned on $H_{t}$
$\bm{\theta}_{0}$ Prior mean of the unknown model parameter $\bm{\theta}_{*}$
$\mathbf{\Sigma}_{0}$ Prior mean of the unknown model parameter $\bm{\theta}_{*}$
$\widehat{\mathbf{\Sigma}}_{t}=\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}\right)^{-1}$ Posterior covariance
$\mathcal{L}_{t}$ Set of labeled points
$\mathcal{U}_{t}$ Set of unlabeled points
$\widehat{\bm{\theta}}_{t,i,j}$ Posterior mean
Table 3: Table of Notations