### v2

# Optimal Design for Adaptive In-Context Prompt Tuning in Large Language Models

Abstract
One emergent ability of large language models (LLMs) is that query-specific examples can be included in the prompt at inference time. In this work, we use active learning for adaptive prompt design and call it Active In-context Prompt Design (AIPD). We design the LLM prompt by adaptively choosing few-shot examples from a training set to optimize performance on a test set. The training examples are initially unlabeled and we obtain the label of the most informative ones, which maximally reduces uncertainty in the LLM prediction. We propose two algorithms, GO and SAL, which differ in how the few-shot examples are chosen. We analyze these algorithms in linear models: first GO and then use its equivalence with SAL. We experiment with many different tasks in small, medium-sized, and large language models; and show that GO and SAL outperform other methods for choosing few-shot examples in the LLM prompt at inference time.
1 Introduction
Large language models (LLMs), such as Vicuna (Chiang et al., 2023), Falcon-40B (Penedo et al., 2023), and OpenLLaMA (Touvron et al., 2023) are applied in mainly two ways: fine-tuning and prompt tuning. In fine-tuning, the LLM weights are adapted to a downstream task (Devlin et al., 2018). Fine-tuning can easily incorporate domain knowledge that a pre-trained model may not possess and resembles classic inductive inference. Fine-tuned models often do not need carefully designed prompts, which makes them easier to deploy. The main drawback of fine-tuning is that it can be costly, because tens of thousands of training examples may be needed to fine-tune billions of parameters of the LLM (Ding et al., 2023). In prompt tuning, the LLM weights are fixed and the LLM is given query-specific examples at inference time that affect its output (Lester et al., 2021). This ability to conduct in-context inference is one of the emergent abilities of LLMs. Prompt tuning does not require large training sets. It is also preferred when query-specific examples are private or change over time, and thus can only be utilized at inference time.
Prior works on prompt tuning mainly focus on hard prompts, which are carefully handcrafted to get the desired output. This can be time-consuming and fragile, as minor prompt modifications can lead to a significant performance drop on the downstream task (Suzgun et al., 2022). In contrast, Zhang et al. (2022a, b) and Diao et al. (2023) explored adaptive prompt design using clustering-based and uncertainty-reducing approaches. While these approaches offer some benefits, we argue that optimal designs (Pukelsheim, 2006; Fedorov, 2013) can outperform them by effectively balancing uncertainty and diversity. Similarly to Zhang et al. (2022a, b) and Diao et al. (2023), we propose a framework for adaptive prompt design called active in-context prompt design (AIPD). The key idea is to design the LLM prompt by adaptively choosing few-shot examples for a set of test examples at inference time. The examples are initially unlabeled and we obtain labels for the most informative ones, which maximally reduce the uncertainty in the LLM prediction for all test examples. We assume that the observed labels are collected from experts (human-in-the-loop) or revealed by an oracle (Dasgupta, 2005; Hanneke et al., 2014). The focus on informativeness and diversity ensures efficient label acquisition by selecting the best examples. This reduces reliance on limited and costly resources such as expert labeling.
[…]
We now state the main contributions of our work:
(1) We propose a G-Optimal design algorithm (GO). The key idea in GO is to retrieve the examples to label that are closest to the set of test examples in the inference task. Our main contribution is the right notion of closeness, based on posterior covariance in a simpler model. GO is implementable with any LLM that can be sampled from, and does not require access to model parameters, feature embeddings of the LLM, or its gradients.
(2) We propose a Simulation-Based Active Learning algorithm (SAL). SAL uses simulation to estimate the impact of labeling unlabeled examples on uncertainty of the example in the inference task. SAL is also implementable with any LLM that can be sampled from.
(3) GO is motivated by optimal designs in linear models (Kiefer and Wolfowitz, 1960; Pukelsheim, 2006). This allows us to analyze GO in linear models (Theorem 1). Our proof is a major departure from similar analyses in fixed-budget best-arm identification in bandits (Azizi et al., 2022; Yang and Tan, 2022), for instance because we directly analyze the discrete allocation problem and each unlabeled example can be labeled at most once. We discuss this in detail right after Theorem 1. SAL is more general than GO because it does not make any linear model assumption in its design. We show that SAL and GO are equivalent in linear models in Theorem 2.
(4) We evaluate GO and SAL on UCI (Markelle Kelly, 1988) and OpenML (Vanschoren et al., 2013) regression and classification tasks, custom NLP datasets, abstract reasoning corpus (ARC) tasks (Alford, 2021; Mirchandani et al., 2023), and Probabilistic Context Free Grammar (PCFG) tasks (Hupkes et al., 2020). GO and SAL consistently outperform other active prompt tuning methods (Zhang et al., 2022a, b; Diao et al., 2023) for choosing few-shot examples in majority of the tasks.
We advance the understanding of active in-context prompt design in LLMs and develop a practical methodology for adaptive prompt design. To our knowledge, this is the first paper that analyzes optimal design based prompting approaches that correctly balance uncertainty and diversity-based sampling as opposed to other existing adaptive prompting-based approaches (Zhang et al., 2022a, b; Diao et al., 2023).
This paper is organized as follows. Section 2 introduces the problem setting. Section 3 presents our methods and discusses their properties. Section 4 is devoted to analyzing our methods. Section 5 validates our approach empirically. We review related work in detail in Appendix A. Finally, Section 6 summarizes our contributions and suggests avenues for future work.
2 Setting
We pose the problem of adaptive prompt design as active learning. We adopt the following standard active learning terminology (Lewis, 1995; Tong and Koller, 2001; Dasgupta, 2005; Dasgupta et al., 2007; Hanneke et al., 2014). We have a $d$-dimensional feature space $\mathcal{X}\subset\mathbb{R}^{d}$ and a $d_{y}$-dimensional label space $\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$. A labeled example is a pair $(\mathbf{x},Y)\in\mathcal{X}\times\mathcal{Y}$. The feature vectors and labels are related as $Y=f(\mathbf{x},\bm{\theta}_{*})+\varepsilon$, where $f$ is an underlying model, $\bm{\theta}_{*}$ is its parameter, and $\varepsilon$ is an independent zero-mean noise vector. Our goal is to learn to estimate $f$ on test examples by labeling training examples. We have a budget $T$ on the maximum number of training examples that can be labeled. This constraint can arise due to multiple reasons. For instance, human labels may be necessary and they are naturally costly. Another reason may be that the machine learning model has a limited capacity for including labeled examples, such as the length of prompts in LLMs (Zhang et al., 2022a, b; Diao et al., 2023).
Now we introduce our notation in detail. Denote $[m]=\{1,2,\ldots,m\}$. We have $n$ training examples $\mathcal{X}_{\text{examples}}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ and $K$ test examples $\mathcal{X}_{*}=\left\{\mathbf{x}_{*,1},\dots,\mathbf{x}_{*,K}\right\}$. We assume that both sets are related, such as being sampled from the same distribution. The label of the test example $\mathbf{x}_{*,k}$ is $Y_{*,k}$. In our motivating theme recognition example the training examples $\mathbf{x}_{i}$ and test example $\mathbf{x}_{*,k}$ is a concatenation of triplets of movies, and the label $Y_{i}$ or $Y_{*,k}$ is the common theme amongst the triplets respectively. We want to infer $Y_{*,k}$ for all $k\in[K]$ without explicitly modeling the complex function $f$. We model the function using an LLM which we treat as a general inference machine because of its large representation capacity (Brown et al., 2020a; Mirchandani et al., 2023). Specifically, let $H_{t}=\{(X_{\ell},Y_{\ell})\}_{\ell\in[t-1]}$ be a set of $t-1$ previously labeled examples, where $X_{\ell}\in\mathcal{X}_{\text{examples}}$ is the $\ell$-th labeled example and $Y_{\ell}$ is its label. Then we denote by $p(\cdot\mid\mathbf{x},H_{t})$ the distribution over labels of an LLM for a queried example $\mathbf{x}$ when $H_{t}$ is used as few-shot in-context examples. To implement this in the LLM, we simply concatenate $\mathbf{x}$ and $H_{t}$ in context (Zhang et al., 2022a, b; Diao et al., 2023). We know that in-context examples affect the distribution of responses of an LLM (Xie et al., 2021; Suzgun et al., 2022; Deng et al., 2023; Lee et al., 2023). So, the problem of learning $f$ under a budget $T$ can be viewed as selecting $H_{T+1}$ such that $p(Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1})$ is high for all test examples $k\in[K]$. This problem is challenging, especially when the training examples need to be labeled.
To effectively reduce the uncertainty of $Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}$, we need to quantify it. One possibility it to use the entropy $-\mathbb{E}_{Y_{*,k}\sim p(\cdot\mid\mathbf{x}_{*,k},H_{T+1})}[\log p(Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1})]$. This is problematic because the entropy is hard to estimate for high-dimensional random variables (Vershynin, 2020), especially without having access to $p(\cdot\mid\mathbf{x}_{*,k},H_{T+1})$ beyond sampling from it. This is a shortcoming of recent adaptive prompting techniques (Zhang et al., 2022a, b; Diao et al., 2023). Therefore, we propose using the covariance of $Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}$ as the uncertainty measure. Specifically, we measure the uncertainty of the $k$-th test example by $\operatorname{tr}(\mathrm{cov}[Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}])$ and the uncertainty over all test examples by $\max_{k\in[K]}\operatorname{tr}(\mathrm{cov}[Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}])$. Since the trace of the covariance is the sum of the variances in individual dimensions, our objective can be interpreted as minimizing the maximum variance over the predicted labels of all test examples. This is a natural measure of uncertainty in linear models and corresponding optimal designs (Pukelsheim, 2006; Fedorov, 2013).
Before we present our algorithms, we wanted to outline their general design. Given a budget $T$, we design sequential adaptive algorithms over $T$ rounds, where the example $X_{t}\in\mathcal{X}_{\text{examples}}$ in round $t\in[T]$ is chosen as a function of $H_{t}=\{(X_{\ell},Y_{\ell})\}_{\ell\in[t-1]}$ up to that round. Since $H_{t}$ summarizes past actions of the algorithm, we call it a history. The label of example $X_{t}$ is $Y_{t}=f(X_{t},\bm{\theta}_{*})+\varepsilon_{t}$, where $\varepsilon_{t}$ is an independent zero-mean noise vector in round $t$. Our objective is to minimize the maximum uncertainty over all test examples, $\max_{k\in[K]}\operatorname{tr}(\mathrm{cov}[Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}])$.
3 Algorithms
Algorithm 1 G-optimal design (GO)
1:  Input: Training set $\mathcal{X}_{\text{examples}}=\{\mathbf{x}_{i}\}_{i=1}^{n}$, test set $\mathcal{X}_{*}=\{\mathbf{x}_{*,k}\}_{k=1}^{K}$, budget $T$
2:  $\mathcal{L}_{1}\leftarrow\emptyset,\ \mathcal{U}_{1}\leftarrow[n],\ H_{1}\leftarrow\{\}$
3:  for $t=1,\dots,T$ do
4:     $I_{t}=\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\left(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)^{-1}\mathbf{x}_{*,k}\,$
5:     $X_{t}\leftarrow\mathbf{x}_{I_{t}}\in\mathcal{X}_{\text{examples}}$
6:     Observe label $Y_{t}$ of example $X_{t}$
7:     $\mathcal{L}_{t+1}\leftarrow\mathcal{L}_{t}\cup\{I_{t}\},\ \mathcal{U}_{t+1}\leftarrow\mathcal{U}_{t}\setminus\{I_{t}\}$
8:     $H_{t+1}\leftarrow H_{t}\cup\{(X_{t},Y_{t})\}$
9:  end for
10:  Output: Sample $Y_{*,k}\sim p(\cdot\mid\mathbf{x}_{*,k},H_{T+1})$ for $k\in[K]$
In this section, we introduce our active learning algorithms for selecting most informative training examples from $\mathcal{X}_{\text{examples}}$. To simplify notation, we assume scalar labels and then discuss an extension to vector labels at the end of the section. We also let $\mathcal{L}_{t}\subseteq[n]$ and $\mathcal{U}_{t}\subseteq[n]$ be the indices of all labeled and unlabeled training examples up to round $t$, respectively. Note that $\mathcal{L}_{t}\cup\mathcal{U}_{t}=[n]$.
3.1 Optimal Design Algorithm
The key idea is to label examples in $\mathcal{X}_{\text{examples}}$ that minimize the maximum uncertainty of predictions over all test examples $\mathbf{x}_{*,k}$. Our computation of uncertainty is borrowed from linear models. Specifically, take a linear model $Y=\mathbf{x}^{\top}\bm{\theta}_{*}+\varepsilon$, where $\bm{\theta}_{*}\in\mathbb{R}^{d}$ is its parameter and $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ is independent noise. Suppose that $\bm{\theta}_{\ast}\sim\mathcal{N}(\bm{\theta}_{0},\mathbf{\Sigma}_{0})$. Then a well-known result in Bayesian statistics (Bishop, 2006) is that the posterior variance of the model estimate at an example $\mathbf{x}_{*,k}$ given labeled examples $H_{t}$ is $\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t}\mathbf{x}_{*,k}$, where $\widehat{\mathbf{\Sigma}}_{t}=(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top})^{-1}$ is the posterior covariance of $\bm{\theta}_{*}\mid H_{t}$. Therefore, the maximum uncertainty over test examples is $\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t}\mathbf{x}_{*,k}$. The key observation is that this quantity does not depend on labels. Therefore, it can be optimized greedily by choosing the training example that minimizes it the most,
$I_{t}=\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\left(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)^{-1}\mathbf{x}_{*,k}\,,$ (1)
where $\mathcal{U}_{t}$ are indices of all unlabeled training examples up to round $t$. After the index $I_{t}$ is chosen, the example $\mathbf{x}_{I_{t}}$ and its label $Y_{t}$ are added to the history to get $H_{t+1}$ for the next iteration $t+1$.
This algorithm is a greedy solution to the G-optimal design (Pukelsheim, 2006; Katz-Samuels et al., 2021). We call it G-Optimal design and abbreviate it as GO. The pseudocode of GO is in Algorithm 1. Note that GO does not depend on observed $Y_{t}$. Similar optimal designs have been effectively applied in active learning (Chaudhuri et al., 2015; Mukherjee et al., 2022), bandits (Fontaine et al., 2021; Mason et al., 2021), and reinforcement learning (Wagenmaker et al., 2022). However, this is the first paper that studies optimal design for adaptively designing prompts (Zhang et al., 2022a; Diao et al., 2023). (1) can be viewed as choosing that training example $\mathbf{x}_{i}\in\mathcal{U}_{t}$ that minimizes the maximum eigenvalue of the posterior covariance $\widehat{\mathbf{\Sigma}}_{t}$. Therefore this leads to reducing the uncertainty over the model parameter $\bm{\theta}_{*}$ as the confidence ellipsoid around $\bm{\theta}_{\star}$ shrinks (Lattimore and Szepesvári, 2020). Note that maximum eigenvalue reduction also ensures diversity as it leads to choosing training examples along all directions in $\mathbb{R}^{d}$.
The time complexity of GO is $O(Kd^{2}nT)$. This is because, for $T$ rounds, the algorithm searches for the best training example out of at most $n$ and evaluates it on all test examples $\mathbf{x}_{*,k}\in\mathcal{X}_{*}$. The evaluation of each test example in round $t$, $\mathbf{x}_{*,k}^{\top}(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}\mathbf{x}_{*,k}$, takes $O(d^{2})$ time, because $(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}$ can be computed in $O(d^{2})$ time using the Sherman-Morrison formula. In the last step, the LLM is queried $K$ times to return $\{Y_{*,k}\}_{k=1}^{K}$.
3.2 Simulation-Based Algorithm
Algorithm 2 Simulation-based active learning (SAL)
1:  Input: Training set $\mathcal{X}_{\text{examples}}=\{\mathbf{x}_{i}\}_{i=1}^{n}$, test set $\mathcal{X}_{*}=\{\mathbf{x}_{*,k}\}_{k=1}^{K}$, budget $T$
2:  $\mathcal{L}_{1}\leftarrow\emptyset,\ \mathcal{U}_{1}\leftarrow[n],\ H_{1}\leftarrow\{\}$
3:  for $t=1,\dots,T$ do
4:     for all $i\in\mathcal{U}_{t}$ do
5:        for all $\mathbf{x}_{*,k}\in\mathcal{X}_{*}$ do
6:           for $j=1,2,\ldots,m$ do
7:              Sample $Y^{(j)}_{t,i}\sim p(\cdot\mid\mathbf{x}_{i},H_{t})$
8:              $H_{t,i,j}\leftarrow H_{t}\cup\{(\mathbf{x}_{i},Y^{(j)}_{t,i})\}$
9:              Sample $\tilde{Y}^{(j,1)}_{t,i,k},\tilde{Y}^{(j,2)}_{t,i,k}\sim p(\cdot\mid\mathbf{x}_{*,k},H_{t,i,j})$
10:           end for
11:        end for
12:     end for
13:     $I_{t}\leftarrow\mathop{\rm arg\,min}_{i\in\mathcal{U}_{t}}\max_{k\in[K]}\frac{1}{m}\sum_{j=1}^{m}\left(\tilde{Y}^{(j,1)}_{t,i,k}-\tilde{Y}^{(j,2)}_{t,i,k}\right)^{2}$
14:     $X_{t}\leftarrow\mathbf{x}_{I_{t}}\in\mathcal{X}_{\text{examples}}$
15:     Observe label $Y_{t}$ of example $X_{t}$
16:     $\mathcal{L}_{t+1}\leftarrow\mathcal{L}_{t}\cup\{I_{t}\},\ \mathcal{U}_{t+1}\leftarrow\mathcal{U}_{t}\setminus\{I_{t}\}$
17:     $H_{t+1}\leftarrow H_{t}\cup\{(X_{t},Y_{t})\}$
18:  end for
19:  Output: Sample $Y_{*,k}\sim p(\cdot\mid\mathbf{x}_{*,k},H_{T+1})$ for $k\in[K]$
While GO reduces uncertainty in label predictions, it has a major limitation. The chosen example $X_{t}$ at round $t$ is not affected by observed labels $(Y_{\ell})_{\ell\in[t-1]}$. This is because (1) does not depend on $(Y_{\ell})_{\ell\in[t-1]}$. While this is a property of linear models, it is undesirable in non-linear models, such as LLMs. To address this limitation, we propose a new algorithm that simulates the impact of labeling examples on the uncertainty of predicted labels. We call it Simulation-Based Active Learning and abbreviated it as SAL. The pseudocode of SAL is provided in Algorithm 2.
The key idea in SAL is to replace the closed-form formula in (1) by a simulation. We detail the algorithm next. Fix round $t$, history $H_{t}$, and candidate example $\mathbf{x}_{i}$. To estimate the impact of labeling $\mathbf{x}_{i}$, we simulate its labels $m$ times. For each simulation $j\in[m]$, we sample $Y^{(j)}_{t,i}$ from the conditional distribution $p(\cdot\mid\mathbf{x}_{i},H_{t})$ using the LLM. Then we extend the history $H_{t}$ by $\mathbf{x}_{i}$ and its simulated label $Y^{(j)}_{t,i}$, $H_{t,i,j}=H_{t}\cup\{(\mathbf{x}_{i},Y^{(j)}_{t,i})\}$. This process results in $m$ copies of augmented histories, each reflecting a potential outcome of labeling of $\mathbf{x}_{i}$. Finally, we take two independent samples for each $j\in[m]$ as $\tilde{Y}^{(j,1)}_{t,i,k},\tilde{Y}^{(j,2)}_{t,i,k}\sim p(\cdot\mid\mathbf{x}_{*,k},H_{t,i,j})$. The maximum uncertainty over test examples after labeling $\mathbf{x}_{i}$ is estimated as
$\max_{k\in[K]}\frac{1}{m}\sum_{j=1}^{m}\left(\tilde{Y}^{(j,1)}_{t,i,k}-\tilde{Y}^{(j,2)}_{t,i,k}\right)^{2}\,.$ (2)
The training example with the lowest value is chosen and we denote its index by $I_{t}$. Then $\mathbf{x}_{I_{t}}$ and its observed label $Y_{t}$ are added to the history to get $H_{t+1}$ for the next iteration $t+1$.
Next we justify SAL. Consider the same setting as in Section 3.1. Given a label $Y^{(j)}_{t,i}$ for example $\mathbf{x}_{i}$, the posterior distribution of $\bm{\theta}_{*}\mid H_{t,i,j}$ is $\mathcal{N}(\widehat{\bm{\theta}}_{t,i,j},\widehat{\mathbf{\Sigma}}_{t,i})$, where $\widehat{\mathbf{\Sigma}}_{t,i}=(\widehat{\mathbf{\Sigma}}_{t}^{-1}+\sigma^{-2}\mathbf{x}_{i}\mathbf{x}_{i}^{\top})^{-1}$ is the simulated posterior covariance of $\bm{\theta}_{*}$ and
$\widehat{\bm{\theta}}_{t,i,j}=\widehat{\mathbf{\Sigma}}_{t,i}\left(\widehat{\mathbf{\Sigma}}_{0}^{-1}\bm{\theta}_{0}+\sigma^{-2}\left(\sum_{\ell=1}^{t-1}X_{\ell}Y_{\ell}+\mathbf{x}_{i}Y^{(j)}_{t,i}\right)\right)$
[…]
The time complexity of SAL is $O(nKmT)$. This is because it searches for the best example out of at most $n$ in $T$ rounds for each test example $k\in[K]$. The evaluation of impact on each test example requires $2m$ LLM queries.
Vector labels: GO and SAL are easy to extend to vector labels, $d_{y}>1$. GO does not depend on labels at all. The only modification in SAL is that (2) is replaced with $\max_{k\in[K]}\frac{1}{m}\sum_{j=1}^{m}\|\tilde{Y}^{(j,1)}_{t,i,k}-\tilde{Y}^{(j,2)}_{t,i,k})\|_{2}^{2}$. This is the sum of the posterior variances of the labels over all dimensions.
4 Analysis
In this section, we analyze GO and SAL. The analysis is under the assumption that the labels are scalar and hence, our objective simplifies to minimizing $\max_{k\in[K]}\mathrm{var}[Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}]$. The analysis is organized as follows. First, we prove that our objective is decreasing in history but not supermodular, which precludes a straightforward analysis. This property of our objective function is proved in Section B.1 and Section B.2. Second, we analyze GO using the closed form of the posterior covariance $\widehat{\mathbf{\Sigma}}_{t}$. Finally, we prove the equivalence of GO and SAL, and thereby provide guarantees for SAL. All analyses are under the assumption of a linear model with Gaussian noise. These proofs are in Section B.3 and Section B.4.
4.1 Analysis of GO
To address challenge posed due to $f$ not being a supermodular (Lemma 5), we leverage the properties of the rank-$1$ updates in GO. The proof is under the assumption that at round $t$, the training examples can be partitioned as $\mathcal{X}=S_{k}\cup\overline{S}_{k}$. The set $S_{k}$ represents examples that are close to $\mathbf{x}_{*,k}$. The set $S_{k}$ is convex such that for a $\alpha_{k}\geq 0$ we have $\mathbf{x}^{\top}\mathbf{y}\geq\alpha_{k}$ for all $\mathbf{x},\mathbf{y}\in S_{k}$. In essence, $\alpha_{k}$ governs the minimum level of similarity required for examples within $S_{K}$ to be considered similar to the test example $\mathbf{x}_{*,k}$. This is achieved by setting a lower bound on the inner product between any two examples in the set. The set $\overline{S}_{k}$ represents examples that are not close to $\mathbf{x}_{*,k}$. It is defined $\beta_{k}\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\leq\beta_{k}$ for all $\mathbf{x}\in S_{k}$ and $\mathbf{y}\in\overline{S}_{k}$. In contrast to $\alpha_{k}$, $\beta_{k}$ limits the maximum similarity any example in $S_{k}$ can have with examples outside this set. Define $\alpha_{\min}=\min_{k}\alpha_{k}$, and $\beta_{\max}=\beta_{\max}$. Define the set $S=\cap_{k=1}^{K}S_{k}$ as the set of all examples that are close to all $\{\mathbf{x}_{*,k}\}_{k=1}^{K}$ and $\overline{S}=\cup_{k=1}^{K}\overline{S}_{k}$ as the set of all examples that are not close to all $\{\mathbf{x}_{*,k}\}_{k=1}^{K}$. Assume $S\neq\{\emptyset\}$ and $|S|>T$. With this in hand, we prove the following claim.
Theorem 1.
Let $\alpha_{\min},\beta_{\max}\geq 0$ be set such that $\beta_{\max}\geq 1-\alpha_{\min}^{2}$ and $T\leq\tfrac{\alpha_{\min}^{2}}{(\beta_{\max}+\sqrt{2})\beta_{\max}d}$. Then for any $\mathbf{x}_{*,k}$ we can show that $\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{T+1}\mathbf{x}_{*,k}\leq\frac{1}{\alpha^{2}_{\max}T+1}+(1-\alpha^{2}_{\max})\,.$
The proof is in Section B.3. It is a major departure from similar proofs in active learning with a fixed budget (Tong and Koller, 2001; Hanneke et al., 2014; Azizi et al., 2022; Yang and Tan, 2022) in three aspects. First, we analyze the discrete allocation problem in (3) instead of its continuous optimal-design relaxation (Pukelsheim, 2006). Second, any unlabeled example in $\mathcal{X}$ is labeled at most once. Finally, (3) is asymmetric in the sense that we optimize the uncertainty of a single example $\mathbf{x}_{*}$ over a larger set. To make the analysis manageable, we impose structure on $\mathcal{X}$. The claim in Theorem 1 holds for any $T$ if $\beta_{\max}=1/(4dn)$ and $\alpha_{\min}=\sqrt{1-\beta_{\max}}$. In this case, $\alpha_{\min}$ is close to $1$, and we get a near-optimal $O(1/T)$ decrease in posterior variance.
4.2 Analysis of SAL
For a sufficiently large sample size $m$ in SAL, we can establish the following equivalence of SAL and GO.
Theorem 2.
Fix a failure probability $\delta\in(0,1)$. Define $\sigma_{t,i,k}^{2}\!\!=\!\!\mathbb{E}[\frac{1}{m}\sum_{j=1}^{m}(\widetilde{Y}^{(j,1)}_{t,i,k}-\widetilde{Y}^{(j,2)}_{t,i,k})^{2}]=2\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\sigma^{2}$, and define $\sigma_{t,i,\max}^{2}=\max_{k\in[K]}\sigma_{t,i,k}^{2}$. Then for any $t\in[T]$ and $i\in\mathcal{U}_{t}$, we have that
$\sigma_{t,i,\max}^{2}\left[1-2\sqrt{\frac{\log(1/\delta)}{m}}\right]$ $\leq\max_{k\in[K]}\frac{1}{m}\sum_{j=1}^{m}\left(\widetilde{Y}^{(j,1)}_{t,i,k}-\widetilde{Y}^{(j,2)}_{t,i,k}\right)^{2}\leq\sigma_{t,i,\max}^{2}\left[1+2\sqrt{\frac{\log(1/\delta)}{m}}+\frac{2\log(1/\delta)}{m}\right]\,.$
Moreover, for $m\geq 8\log(1/\delta)$ we have that
$2\max_{k}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\frac{\sigma^{2}}{2}\leq\max_{k}\frac{1}{m}$ $\sum_{j=1}^{m}\left(\widetilde{Y}^{(j,1)}_{t,i,k}-\widetilde{Y}^{(j,2)}_{t,i,k}\right)^{2}\leq 5\max_{k}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\frac{5\sigma^{2}}{2}.$
These claims hold with probability at least $1-\delta$.
The claim is proved in Section B.4. The key idea in the proof is that (2) multiplied by $m/[2(\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\sigma^{2})]$ is a chi-squared random variable with $m$ degrees of freedom. Then we use concentration inequalities of Laurent and Massart (2000) to get a high-probability confidence interval on distance to the mean $m$, which in turn allows us to relate the actual variance to its empirical estimate. Theorem 2 shows that SAL is equivalent to GO for a sufficiently large sample size $m$. Theorem 1 can be then adapted to SAL as follows. The only change is in condition on $T$, which changes to $T\leq\alpha^{2}\bigg{/}\left(\beta+\sqrt{2}O\left(\tfrac{1-\sqrt{1/m}}{1+\sqrt{1/m}}\right)\right)\beta d+O(\sqrt{1/m})\,.$ Therefore, SAL attains a near-optimal $O(1/T)$ decrease in posterior variance as $m\to\infty$.
5 Experiments
We evaluate GO and SAL on a variety of prediction tasks. These tasks cover both classification and regression, including natural language features, and help us to evaluate the capabilities of GO and SAL to choose few-shot examples for active in-context prompt design. We also demonstrate that GO and SAL can be used for general pattern recognition. Detailed descriptions of all datasets are in Section C.1. We describe the prompts in detail in Appendix D.
5.1 Experimental Setup
We use Mistral-7B (Jiang et al., 2023), Vicuna-13B (Chiang et al., 2023), and Falcon-40B (Penedo et al., 2023) as the LLMs and design prompts following Dinh et al. (2022) and Suzgun et al. (2022). To investigate the impact of LLM model size on performance, we experiment with these three models of varying sizes: 7B, 13B, and 40B. Interestingly, we observe that the smaller models (Mistral-7B and Vicuna-13B) perform very poorly on certain tasks. Examples of the prompts are given in Appendix D. Each experiment is averaged over $10$ trials. At the beginning of each trial, we randomly select $K=20$ test examples. We describe in detail how the training set and $n$ are chosen for each dataset in Appendix C.
Each run is a simulation that proceeds as follows. In round $t$, each method selects a training example to label $X_{t}$ and then observes the true label $Y_{t}$. All past interactions $H_{t}=\left\{(X_{\ell},Y_{\ell})\right\}_{\ell\in[t-1]}$ along with the test examples $\mathbf{x}_{*,k}$ are used to craft a prompt for the LLM. The performance at round $t$ is evaluated by the error $\mathcal{L}(t)=\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}(Y_{*,k},\tilde{Y}_{*,k,t})$, where $Y_{*,k}$ is the true label of test example $\mathbf{x}_{*,k}$, $\tilde{Y}_{*,k,t}$ is its LLM predicted label in round $t$, and $\mathcal{L}(y_{*},y)$ is a task-specific error function. For classification tasks, we choose $\mathcal{L}(y_{*},y)=\mathbb{I}\{y_{*}=y\}$ and call $\mathcal{L}(t)$ a misclassification error. For regression tasks, we choose $\mathcal{L}(y_{*},y)=(y_{*}-y)^{2}$ and call $\mathcal{L}(t)$ the MSE. For pattern recognition tasks, where $Y_{*,k}$ and $\tilde{Y}_{*,k,t}$ are either vectors or matrices, we choose let $\mathcal{L}(y_{*},y)=\mathbb{I}\{y_{*}=y\}$ and $\mathcal{L}(t)$ represents 0-1 error.
We posit that GO and SAL perform well because they both reduce the uncertainty of test examples based on the right notion of similarity. To show this, we compare to baselines that reduce uncertainty uniformly (like Uniform), or reduce uncertainty informatively (Least or Max-Entropy), or only select examples with similar features to test examples (Greedy-NN). As shown in our extensive experiments, these baselines fail to match the capabilities of GO and SAL to select informative examples in the majority of the tasks. The following methods are compared in our experiments:
(1) Uniform: The example $X_{t}$ in round $t$ is sampled uniformly at random from the unlabeled set $\mathcal{U}_{t}$. Uniform is a pure exploration algorithm that does not take into account the similarity to test examples and variance reduction. We chose it as a baseline because it tends to work well in practice. Therefore, it is used frequently in active learning and prompt tuning papers (Zhang et al., 2022b; Diao et al., 2023).
(2) Greedy-NN: The example $X_{t}$ in round $t$ is chosen to align the most with all test examples $\mathbf{x}_{*,k}$ such that $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\mathbf{x}_{i}$. This baseline shows that our information gathering rule in (1) goes beyond pure feature similarity. This baseline is similar to the automatic exemplar construction method by clustering by Zhang et al. (2022b).
(3) Least: This is similar to the disagreement-based method of Diao et al. (2023). The disagreement score of the example $i\in\mathcal{U}_{t}$ is calculated as $s_{i}=\sum_{k=1}^{K}Y_{tik}$ where $Y_{tik}\sim p\left(\cdot\mid\mathbf{x}_{*,k},\mathbf{x}_{i}\right)$ is the number of unique answers by for test example $\mathbf{x}_{*,k}$ using only $\mathbf{x}_{i}$ as the in-context example by the LLM. Then the example selected at round $t$ is $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}s_{i}$. This is the unlabeled example where the LLM, disagrees the most for all test examples and is least confident. We compare against this baseline to show that our information gathering rule in (1) goes beyond just uncertainty sampling but also takes into account the diversity of training examples when choosing to label the next example.
(4) Max-Entropy: This is the uncertainty-based maximum entropy method of Zhang et al. (2022a); Diao et al. (2023). At round $t$ the example with the highest entropy is selected as $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}-\sum_{k=1}^{K}\bar{Y}_{tik}\ln\bar{Y}_{tik}$ where $\bar{Y}_{tik}\sim p\left(\cdot\mid\mathbf{x}_{*,k},\mathbf{x}_{i}\right)$ is the frequency of a predicted answer among all predictions for the test example $\mathbf{x}_{*,k}$ using $\mathbf{x}_{i}$ as the in-context example by the LLM. A larger entropy denotes greater uncertainty and therefore, an unlabeled example with the largest entropy will be selected. Again we compare against this uncertainty-based baseline to show that our information gathering rule in (1) goes beyond just uncertainty sampling but also considers the diversity of training examples when choosing to label the next example.
(5) GO (ours): This is Algorithm 1 where $\mathcal{X}$ are the original feature vectors.
(6) GO-Inst (ours): This is Algorithm 1 where the original feature vectors are used in the prompt but $\mathcal{X}$ are their $768$-dimensional Instructor embeddings (Su et al., 2022). We use this for natural language classification tasks.
(7) SAL (ours): This is Algorithm 2 where $\mathcal{X}$ are the original feature vectors. To implement SAL efficiently, we combine it with GO as a preprocessing step. Specifically, in round $t$, GO first chooses $5$ most informative examples from $\mathcal{U}_{t}$ and then we apply SAL. We use $m=1$ in all experiments. We use these approximations because SAL is computationally expensive (Section 3.2). Similarly to GO-Inst, we use Instructor embeddings for natural language classification tasks.
Datasets Uniform Greedy-NN Least Max-Entropy GO (ours) SAL (ours) iris $0.41\pm 0.11$ $0.60\pm 0.13$ $0.64\pm 0.15$ $0.72\pm 0.17$ $0.38\pm 0.14$ $\mathbf{0.34\pm 0.14}$ M banknote $0.75\pm 0.10$ $\mathbf{0.58\pm 0.04}$ $0.59\pm 0.02$ $0.73\pm 0.16$ $0.77\pm 0.07$ $0.75\pm 0.15$ balance-scale $0.61\pm 0.13$ $0.69\pm 0.22$ $0.55\pm 0.25$ $0.57\pm 0.14$ $\mathbf{0.48\pm 0.09}$ $0.72\pm 0.04$ thyroid-new $\mathbf{0.44\pm 0.12}$ $0.70\pm 0.08$ $0.74\pm 0.12$ $0.57\pm 0.08$ $0.55\pm 0.06$ $0.63\pm 0.14$ iris $0.22\pm 0.24$ $0.60\pm 0.37$ $0.60\pm 0.49$ $0.40\pm 0.20$ $0.20\pm 0.24$ $\mathbf{0.20\pm 0.24}$ V banknote $0.40\pm 0.37$ $0.80\pm 0.24$ $0.50\pm 0.32$ $0.50\pm 0.32$ $0.50\pm 0.32$ $\mathbf{0.10\pm 0.20}$ balance-scale $0.60\pm 0.20$ $0.60\pm 0.37$ $0.50\pm 0.32$ $0.80\pm 0.24$ $\mathbf{0.30\pm 0.24}$ $0.50\pm 0.00$ thyroid-new $0.52\pm 0.45$ $1.00\pm 0.00$ $0.70\pm 0.24$ $0.50\pm 0.00$ $0.60\pm 0.20$ $\mathbf{0.50\pm 0.32}$ iris $\mathbf{0.20\pm 0.06}$ $0.62\pm 0.14$ $0.70\pm 0.20$ $0.65\pm 0.18$ $0.42\pm 0.10$ $0.33\pm 0.23$ F banknote $0.45\pm 0.23$ $0.53\pm 0.25$ $0.60\pm 0.12$ $0.42\pm 0.17$ $0.45\pm 0.06$ $\mathbf{0.45\pm 0.1}$ balance-scale $0.70\pm 0.28$ $0.68\pm 0.13$ $0.85\pm 0.12$ $0.62\pm 0.08$ $0.47\pm 0.24$ $\mathbf{0.45\pm 0.13}$ thyroid-new $0.55\pm 0.29$ $0.57\pm 0.20$ $0.75\pm 0.19$ $0.65\pm 0.15$ $0.55\pm 0.23$ $\mathbf{0.53\pm 0.12}$
Table 1: Misclassification error in classification datasets using Mistral-7B (M), Vicuna-13B (V), and Falcon-40B (F) on $K=20$ test examples at the end of budget $T=5$.
Datasets Uniform Greedy-NN Least Max-Entropy GO (ours) SAL (ours) machine(e+04) $11.4\pm 3.34$ $10.5\pm 2.44$ $14.3\pm 3.39$ $11.0\pm 1.93$ $\mathbf{10.5\pm 3.74}$ $10.6\pm 3.84$ M fifa(e-04) $1.40\pm.216$ $3.72\pm 1.12$ $1.18\pm.53$ $4.15\pm 1.11$$9$404 .$\mathbf{.68\pm.26}$ machine(e+04) $5.59\pm 1.35$ $5.04\pm.851$ $7.95\pm 1.69$ $4.98\pm 1.06$ $5.66\pm 1.54$ $\mathbf{4.80\pm 1.46}$ V fifa(e+03) $5.90\pm 1.59$ $4.72\pm.931$ $5.11\pm 1.12$ $6.76\pm 1.69$ $\mathbf{1.44\pm.258}$ $2.59\pm.742$ machine(e+03) $1.16\pm 1.22$ $4.28\pm 2.30$ $2.15\pm 1.08$ $033.50\pm 2.45e+$ 03 .$32\pm.209$ $\mathbf{2.96\pm 1.56}$ F fifa(e+01) $7.78\pm 3.85$ $6.95\pm 2.93$ $12.4\pm 12.3$ $26.3\pm 37.6$ $7.90\pm 4.64$ $\mathbf{4.61\pm 4.35}$
Table 2: MSE in regression datasets using Mistral-7B (M), Vicuna-13B (V), and Falcon-40B (F) on $K=20$ test examples at the end of budget $T=5$.
Task Uniform Greedy-NN Least Max-Entropy GO (ours) SAL (ours) Arc-1 $0.45\pm 0.50$ $0.45\pm 0.50$ $0.90\pm 0.30$ $0.60\pm 0.49$ $0.30\pm 0.46$ $\mathbf{0.15\pm 0.36}$ Arc-2 $0.80\pm 0.40$ $1.00\pm 0.00$ $0.80\pm 0.40$ $0.80\pm 0.40$ $0.80\pm 0.40$ $\mathbf{0.01\pm 0.01}$ PCFG-1 $0.60\pm 0.49$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.01$ $0.20\pm 0.40$ $\mathbf{0.02\pm 0.01}$ PCFG-2 $0.20\pm 0.40$ $1.00\pm 0.00$ $0.20\pm 0.40$ $1.00\pm 0.00$ $0.20\pm 0.40$ $\mathbf{0.14\pm 0.40}$
Table 3: 0-1 error using Falcon-40B on $K=20$ test examples at the end of budget $T=5$. ARC-1 is the expansion-contraction task, ARC-2 is the rotation task, PCFG-1 is the add-subtract task, and PCFG-2 is the repeat experiment task. Mistral-7B and Vicuna-13B perform very poorly on these tasks and thus are omitted.
Datasets Uniform Greedy-NN Least Max-Entropy GO (ours) SAL (ours) movie $0.32\pm 0.17$ $0.90\pm 0.06$ $0.87\pm 0.09$ $0.55\pm 0.18$ $\mathbf{0.27\pm 0.10}$ $0.49\pm 0.11$ M entity $0.69\pm 0.19$ $0.86\pm 0.06$ $0.87\pm 0.09$ $0.59\pm 0.15$ $0.65\pm 0.18$ $\mathbf{0.39\pm 0.19}$ theme $0.74\pm 0.05$ $0.74\pm 0.09$ $0.82\pm 0.16$ $\mathbf{0.68\pm 0.09}$ $0.80\pm 0.09$ $0.81\pm 0.08$ movie $0.10\pm 0.20$ $0.70\pm 0.24$ $0.90\pm 0.20$ $0.30\pm 0.24$ $\mathbf{0.02\pm 0.01}$ $0.10\pm 0.20$ V entity $0.20\pm 0.24$ $0.90\pm 0.20$ $0.70\pm 0.24$ $0.60\pm 0.20$ $\mathbf{0.10\pm 0.20}$ $0.10\pm 0.20$ theme $0.90\pm 0.20$ $0.70\pm 0.40$ $1.00\pm 0.00$ $0.70\pm 0.24$ $\mathbf{0.60\pm 0.37}$ $0.80\pm 0.40$ movie $0.55\pm 0.06$ $0.62\pm 0.18$ $0.78\pm 0.17$ $0.53\pm 0.22$ $\mathbf{0.38\pm 0.18}$ $0.47\pm 0.17$ F entity $0.55\pm 0.23$ $0.62\pm 0.08$ $0.75\pm 0.14$ $0.65\pm 0.24$ $0.47\pm 0.23$ $\mathbf{0.42\pm 0.24}$ theme $0.68\pm 0.20$ $0.70\pm 0.13$ $0.85\pm 0.05$ $0.85\pm 0.09$ $\mathbf{0.53\pm 0.18}$ $0.55\pm 0.19$
Table 4: Misclassification error in natural language classification tasks using Mistral-7B (M), Vicuna-13B (V), and Falcon-40B (F) on $K=20$ test examples at the end of budget $T=5$.
All used datasets and experimental setups are described in Appendix C. This section only summarizes the main results.
Standard classification and regression tasks. We use $4$ classification and $2$ regression datasets from UCI and OpenML (Section C.1). We set $T=5$ to simulate the realistic scenario when the test queries provided by the user need to be inferred quickly. For classification tasks, $K=20$ test examples are chosen among the different classes of the dataset. We describe in detail how the training set and $n$ are chosen for each dataset in Appendix C. For regression tasks, $K=20$ random test examples are chosen. Our results on classification tasks are reported in Table 1 and on regression tasks in Table 2. We observe that GO and SAL are the best-performing methods in the majority of the datasets. Note that there is no single baseline that consistently outperforms them.
General pattern recognition. We experiment with $4$ tasks: ARC expansion and contraction, ARC rotation, PCFG Add-Subtract, and PCFG Repeat. Both inputs and outputs in these tasks are vectors or matrices. We describe examples of ARC and PCFG tasks in detail in Section C.1. Each dataset comprises examples of two patterns: expansion and contraction, clockwise and counter-clockwise rotation, add and subtract, repeat first and second digits. In each trial, we choose $K=20$ different test examples equally from two patterns and set $T=5$. Our results are reported in Table 3. In all tasks, GO and SAL are the best-performing methods. SAL outperforms GO in ARC (Alford, 2021; Mirchandani et al., 2023) and PCFG (Hupkes et al., 2020) consistently.
Natural language classification tasks (NLC). We show that GO and SAL work well on general NLP tasks where no explicit numerical features are available. We create three synthetic datasets based on the following tasks: (i) movie-names: predicting a genre from a movie name (e.g., romance, horror), (ii) movie-theme: predicting a common theme for a pair of movie names (e.g., coming-of-age, sci-fi), and (iii) entity-names: predicting an entity’s type from its name (e.g., celebrity, mountain, river). Each dataset comprises $5$ classes. Further details regarding the additional datasets are provided in Appendix C. In each trial, $K=20$ test examples were randomly chosen across the five classes. The feature vectors in GO and SAL are Instructor embeddings of the original text features. Our results are reported in Table 4. We observe again that GO and SAL are the best-performing methods in the majority of the datasets. There is no single baseline that consistently outperforms them. This shows that the optimal design-based approach of GO and SAL correctly balances uncertainty and diversity-based sampling.
6 Conclusions
In this paper, we studied the framework of active in-context prompt design (AIPD) that uses optimal design to systematically choose the most informative unlabeled examples to label for a set of test examples. These informative labeled examples are then used to minimize the prediction error of the LLM for all the test examples. To our knowledge, this is the first paper that studies optimal design for adaptive prompt design. Inspired by the linear model, we proposed an algorithm GO that strategically chooses the most informative examples that minimize the variance of the posterior covariance for any test example from the test set. We proposed a second algorithm SAL that uses simulations to estimate the impact of how unlabeled examples reduce LLMuncertainty for all test examples. It then chooses to label examples that maximally reduce the uncertainty of the LLM for all test examples from the test set. We theoretically analyze GO and SAL and show their equivalence in linear models. We show that both algorithms guarantee information gain at each iteration. Finally, we show empirically that, when used with LLMs like Mistral-7B, Vicuna-13B, and Falcon-40B, both GO and SAL result in better prediction accuracy than other baselines (Zhang et al., 2022a, b; Diao et al., 2023) on tasks like classification, regression, ARC, PCFG, and natural language generation. Our research opens up exciting new directions for future work such as extending AIPD framework beyond text to enable informative example selection for tasks involving images, videos, or other modalities using multi-modal LLMs (Yin et al., 2023). Additionally, the integration of active learning with diffusion models, a powerful class of generative models, presents promising directions for future research (Ho et al., 2020).
Appendix A Related Work
We study the problem of choosing human demonstrations adaptively to get the desired output from the LLM as quickly as statistically possible. We use active learning to choose them and then ask a human to label them. Finally, the human demonstrations are fed as in-context input to the LLM together with the main user query, to obtain the desired output. Thus the name is active transductive inference. Prior works on prompt-tuning and transductive inference [Lester et al., 2021, Dong et al., 2022, Zhang et al., 2022a, Min et al., 2022, Wu et al., 2022, Yu et al., 2022, Suzgun et al., 2022, Liu et al., 2023a, Yu et al., 2023, Liu et al., 2023b] focus on hard prompt-tuning where the user must carefully handcraft the prompt to get the desired output for tasks like movie recommendation, disambiguation QA, navigation, etc. Such examples of carefully handcrafting hard prompts with demonstrations can be found in Suzgun et al. [2022], Srivastava et al. [2022]. These papers also show how failing to design such prompts with demonstrations can lead the LLM to predict wrong outputs. Note that we adaptively design the prompt through carefully chosen demonstrations. In our experiments, we show significant improvement over randomly chosen demonstrations.
The problem of active learning is also related to dataset augmentation [Dukler et al., 2021]. In this work, the most informative unlabeled examples are chosen by optimizing the error of the model on the validation set. The gradient of the validation set error with respect to the weights of unlabeled examples has a closed form when the original model is linearized. The main difference in active learning, including in our work, is that labeling all unlabeled examples would be costly. Therefore, the labels are not available in advance and are queried adaptively. We also learn in context and do not assume that the gradient information of the LLM is available. Prompt composition has also been an active area of research. Bowman et al. [2023] proposed a-la-carte prompt tuning, where prompts are tuned on individual datasets and composed at inference time to mimic the performance of the model that would have been trained on the union of the corresponding datasets. This idea has been further extended by Perera et al. [2023], where prompts for previously unseen tasks are obtained by linearly combining prompts from known tasks. To do this, they use spectral decomposition and project prompts from known tasks to a lower dimensional space. In our work, we do not tune prompts or compose simpler models. We actively probe an LLM, treated as a black box without any extra side information, to answer a test example as accurately as possible, with as little variance in the answer as possible.
Recently there has been a lot of progress in prompt tuning (or aligning). Hassan et al. [2023] studies prompt aligning for a single test sample to adapt to multi-modal test prompts at test time by minimizing the feature distribution shift to the test domain. In contrast in this paper, we study adapting prompts for many test samples without the feature distribution shift assumption. The Wang et al. [2023a] trains a smaller LLMto select demonstrations for a larger LLM. However, we rely on active learning to select the smallest number of informative prompts to be labeled by human labelers. This avoids finetuning a smaller LLM for individual tasks. Wang et al. [2023c] studies transductive inference for diffusion models for a different setting where given a pair of task-specific example images, such as depth from/to image and scribble from/to image, and text guidance, the model automatically understands the underlying task and performs the same task on a new query image following the text guidance. However, in our setting, we do not explicitly encode any guidance text. The Wen et al. [2023] proposes to mix the nearest neighbor method with gradient optimization to select prompts during test time. Similarly, Zhang et al. [2023] proposes a nearest neighbor approach to select in-context examples for computer vision tasks. We compare our approach against such nearest-neighbor selection algorithms. Finally, Bai et al. [2023], Wang et al. [2023b] analyze transductive inference theoretically to understand its universality, generalization capability, and limitations. In contrast, we only do a theoretical analysis of AIPD to show it maximally reduces the estimated variance of the answer to the user’s query. The Zhang et al. [2022b] study the chain-of-thought prompting where they automatically select the prompt using a clustering-based approach.
There are some related works in the area of medical diagnosis chatbot examples that we shared in the introduction. One notable study Caruccio et al. [2024], although not utilizing machine learning techniques, provides valuable insights with its implementation of three hardcoded prompt designs for accurate diagnosis. These designs, however, do not engage in active learning, as they lack the capability to adapt based on user input. Similarly, Kuroiwa et al. [2023] gives more insights into self-diagnostics of orthopedic diseases using ChatGPT. In contrast, web applications such as Buoy Health BuoyHealth and Live Healthily livehealthily employ a more dynamic approach, actively tailoring subsequent questions based on users’ responses. This aligns with active learning principles but it is not clear what techniques they apply and is notably underexplored in academic literature, indicating a potential area for further research. Our setting also goes beyond the single shot active prompt tuning studied in Margatina et al. [2023]. Note that this work studies prompt tuning only for one iteration, and does not take into account the historical context. So it has limited ability for complex tasks like ARC [Alford, 2021] and PCFG [Hupkes et al., 2020] as well as handling vector labels like GO and SAL.
Active Learning (AL): Recently, there has been a lot of focus on using deep AL to finetune LLMs. All AL algorithms tend to balance uncertainty and diversity in the selection of unlabeled examples. We briefly discuss them below and also highlight the main difference of these approaches with prompt aligning with GO and SAL.
(1) Coreset: This is a pure diversity-based approach using a coreset selection. In every iteration, first, the embedding of each unlabeled example is computed from the network’s penultimate layer, and then unlabeled examples are selected using a greedy furthest-first traversal conditioned on all labeled examples [Sener and Savarese, 2017, Geifman and El-Yaniv, 2017, Citovsky et al., 2021]. Observe that in our setting we do not have access to the penultimate layer of the LLM.
(2) Least: This is an uncertainty-based active learning algorithm. Here, the uncertainty score of an unlabeled example is its predicted class probability. At every iteration, this algorithm then samples unlabeled examples with the smallest uncertainty scores [Settles, 2009, 2011, Wang and Shang, 2014].
(3) Margin: This is also an uncertainty-based active learning algorithm [Tong and Koller, 2001, Balcan et al., 2009, Settles, 2009]. At every iteration $t$ it selects unlabeled examples that are sorted according to their multiclass margin score and then selects unlabeled examples that are the hardest to discriminate and can be thought of as examples closest to their class margin. However, in our setting, we do not have any information on the hypothesis space of the LLM and hence cannot implement such a baseline.
(4) Entropy: This is also an uncertainty-based active learning algorithm Wang and Shang [2014], Kremer et al. [2014], Diao et al. [2023]. At every iteration $t$ it selects unlabeled examples according to the entropy of the example’s predictive class probability distribution. We show that Greedy-NN-max-mean is outperformed significantly by GO and SAL in the prediction, pcfg, or arc tasks.
(5) Badge: This is an algorithm that combines both uncertainty and diversity sampling [Ash et al., 2019, 2021]. For each unlabeled example $\mathbf{x}$ its gradient embedding $g_{\mathbf{x}}$ is computed with respect to the parameters of the model’s penultimate layer. Finally, Badge chooses a batch of samples to sample by applying $k$-Means++ [Arthur and Vassilvitskii, 2006] on the gradient embeddings. Again recall that we cannot implement such a baseline as we do not have access to the LLMs last layer.
(6) Badge-KM: This algorithm is similar to Badge but in the final step instead of $k$-Means++ it uses $k$-Means on the gradient embeddings. In Yuan et al. [2020] it is observed that applying $k$-Means on the embeddings results in an increase in accuracy over baselines in some datasets. Further Yuan et al. [2020] observed from the t-SNE plots that $k$-Means select centers that are further apart compared to the ones chosen by $k$-Means++ which leads to more diverse sampling in batches.
(7) Bald: Bayesian Active Learning by Disagreements [Kirsch et al., 2019, Gal et al., 2017] chooses unlabeled examples that are expected to maximize the information gained from the model parameters $\bm{\theta}_{t}$, i.e. the mutual information between predictions and model posterior.
A more comprehensive survey on how AL is used for finetuning deep models can be found in Ren et al. [2021], Zhan et al. [2022]. The Bhatt et al. [2024] study how experimental design can be used to select prompts for finetuning a pre-trained LLM. Some recent works have also focused on selecting unlabeled examples only within a task Wei et al. [2021], Chen et al. [2023], Fifty et al. [2021]. However, these works are geared towards selecting prompts within a task for finetuning, whereas we focus on adaptive prompt design using experimental design. The work of Perlitz et al. [2023] also uses AL for finetuning prompts for LLMsto improve label efficiency. The Kung et al. [2023] proposes an AL framework for instruction tuning. However, their approach again focuses on selecting unlabeled examples inside each task and discriminating one task from another. However, they make the simplifying assumption that all unlabeled examples inside the tasks are equally informative which may inhibit the quality of the selected subset.
Appendix B Proofs
This section contains the properties of our objective and proofs of our main claims.
B.1 Properties of our objective
By the total variance decomposition for a linear model with observation variance $\sigma^{2}$, we get
$\max_{k\in[K]}\mathrm{var}[Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}]$ $=\max_{k\in[K]}\mathrm{var}[\mathbb{E}[Y_{*,k}\mid\mathbf{x}_{*,k},\theta_{*},H_{T+1}]\mid\mathbf{x}_{*,k},H_{T+1}]$
$\quad+\max_{k\in[K]}\mathbb{E}[\mathrm{var}[Y_{*,k}\mid\mathbf{x}_{*,k},\theta_{*},H_{T+1}]\mid\mathbf{x}_{*,k},H_{T+1}]$
$=\max_{k\in[K]}\sigma^{2}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t}^{-1}\mathbf{x}_{*,k}+\sigma^{2}\,.$
Therefore, the variance minimization of $\max_{k\in[K]}Y_{*,k}\mid\mathbf{x}_{*,k},H_{T+1}$ is a combinatorial optimization problem. It can be formulated as follows
$\textstyle S_{*}=\mathop{\rm arg\,min}_{S\subseteq[n],|S|=T}f(S)\,,$ (3)
where the function $f(S)=\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{i\in S}\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)^{-1}\mathbf{x}_{*,k}$ represents the uncertainty associated with any subset $S$, and $S_{*}$ denotes the optimal subset of size $T$. If $f(S)$ was monotone and supermodular in $S$, we would have $(1-1/e)$-optimality guarantees for GO [Nemhauser et al., 1978]. We start with proving the monotonicity of $f(S)$.
Lemma 3.
Function $f(S)$ is monotonically decreasing.
The claim is proved in Section B.2. Now we state the definition of a supermodular function and show that $f(S)$ is not supermodular.
Definition 4 (Supermodular function).
Let $[n]$ be a set of $n$ elements. A function $f:2^{[n]}\to\mathbb{R}$ is supermodular if it satisfies the property of diminishing marginal returns. For any $S_{1}\subseteq S_{2}\subseteq[n]$ and $\mathbf{x}\notin S_{2}$, we have $f(S_{1}\cup\left\{\mathbf{x}\right\})-f(S_{1})\leq f(S_{2}\cup\left\{\mathbf{x}\right\})-f(S_{2})$.
Lemma 5.
The function $f(S)$ is not supermodular.
Proof.
Take $\mathbf{\Sigma}_{0}=I_{d}$, fix $K=1$, $\mathbf{x}_{1}=(1/\sqrt{2},1/\sqrt{2})$, $\mathbf{x}_{2}=(0,1)$, and $\mathbf{x}_{\ast}=(1,0)$. Then $f(\left\{2\right\})-f(\emptyset)=0<0.035714\approx f(\left\{1,2\right\})-f(\left\{1\right\})$. ∎
B.2 Proof of Lemma 3
Recall that the objective function is $f(S)=\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{i\in S}\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)^{-1}\mathbf{x}_{*,k}$. Without loss of generality, let $\sigma^{2}=1$. Fix any $j\in[n]\setminus S$ and let $S_{+}=S\cup\left\{j\right\}$. Then the objective value at $S_{+}$ can be written as
$f(S_{+})=\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\left(\mathbf{\Sigma}_{0}^{-1}+\sum_{i\in S_{+}}\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)^{-1}\mathbf{x}_{*,k}=\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\left(\mathbf{A}+\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\right)^{-1}\mathbf{x}_{*,k}\,,$
where $\mathbf{A}=\mathbf{\Sigma}_{0}^{-1}+\sum_{i\in S}\mathbf{x}_{i}\mathbf{x}_{i}^{\top}$. By the Sherman–Morrison formula, we have
$\left(\mathbf{A}+\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\right)^{-1}=\mathbf{A}^{-1}-\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}\,.$
Now note that $\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}$ is a positive semi-definite matrix for any $\mathbf{x}_{j}$ and any positive semi-definite matrix $\mathbf{A}$. As a result, $\mathbf{x}_{*,k}^{\top}\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}\mathbf{x}_{*,k}\geq 0$ holds for any $\mathbf{x}_{*,k}$ and thus
$f(S_{+})=\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\mathbf{A}^{-1}\mathbf{x}_{*,k}^{\top}-\mathbf{x}_{*,k}^{\top}\frac{\mathbf{A}^{-1}\mathbf{x}_{j}\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}}{1+\mathbf{x}_{j}^{\top}\mathbf{A}^{-1}\mathbf{x}_{j}}\mathbf{x}_{*,k}\leq\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\mathbf{A}^{-1}\mathbf{x}_{*,k}^{\top}=f(S)\,.$
This proves our claim.
B.3 Proof of Theorem 1
The proof is under the assumption that at round $t$, the training examples can be partitioned as $\mathcal{X}_{\text{examples}}=S_{k}\cup\overline{S}_{k}$. The set $S_{k}$ represents examples that are close to $\mathbf{x}_{*,k}$. The set $S_{k}$ is convex and define $\alpha_{k}\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\geq\alpha_{k}$ for all $\mathbf{x},\mathbf{y}\in S_{k}$. The set $\overline{S}_{k}$ represents examples that are not close to $\mathbf{x}_{*,k}$. It is defined $\beta_{k}\geq 0$ such that $\mathbf{x}^{\top}\mathbf{y}\leq\beta_{k}$ for all $\mathbf{x}\in S_{k}$ and $\mathbf{y}\in\overline{S}_{k}$. Define $\alpha_{\min}=\min_{k}\alpha_{k}$, and $\beta_{\max}=\beta_{\max}$.
Define the set $S=\cap_{k=1}^{K}S_{k}$ as the set of all examples that are close to all $\{\mathbf{x}_{*,k}\}_{k=1}^{K}$ and $\overline{S}=\cup_{k=1}^{K}\overline{S}_{k}$ as the set of all examples that are not close to all $\{\mathbf{x}_{*,k}\}_{k=1}^{K}$. Assume $S\neq\{\emptyset\}$ and $|S|>T$.
The inverse of the covariance matrix after $t-1$ observations is
$\widehat{\mathbf{\Sigma}}_{t}^{-1}=\mathbf{\Lambda}_{t}=\mathbf{I}_{d}+\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}=\sum_{i=1}^{d}\lambda_{t,i}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\,.$
The latter is the eigendecompositon of $\mathbf{\Lambda}_{t}$, where $\lambda_{t,i}$ is the $i$-th largest eigenvalue and $\mathbf{v}_{t,i}$ is the corresponding eigenvector. Note that $\mathbf{\Lambda}_{t}^{-1}=\sum_{i=1}^{d}\lambda_{t,i}^{-1}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}$. To simplify exposition, we assume that all examples have unit length. We analyze the eigenvalues of $\mathbf{\Lambda}_{t}$ first.
Lemma 6.
For all $i\in[d]$, $1\leq\lambda_{t,i}\leq t$. Moreover, let $X_{\ell}\in S$ hold for all $\ell\in[t-1]$. Then $\lambda_{t,1}\geq\alpha_{\min}^{2}(t-1)+1$.
Proof.
The first claim follows directly from the definition of $\mathbf{\Lambda}_{t}$ and that $\|X_{\ell}\|_{2}\leq 1$. The second claim is proved using the definition of the maximum eigenvalue,
$\lambda_{t,1}=\mathbf{v}_{t,1}^{\top}\mathbf{\Lambda}_{t}\mathbf{v}_{t,1}\geq\mathbf{x}_{*,k}^{\top}\left(\mathbf{I}_{d}+\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}\right)\mathbf{x}_{*,k}\geq\alpha_{\min}^{2}(t-1)+1\,.$
The last inequality follows from $\mathbf{x}_{*,k}^{\top}X_{\ell}\geq\alpha_{\min}$. This completes the proof. ∎
We continue with claims about the eigenvectors of $\mathbf{\Lambda}_{t}$.
Lemma 7.
Let $X_{\ell}\in S$ hold for all $\ell\in[t-1]$. Then $\mathbf{v}_{t,1}\in S$. Moreover, let $\beta_{\max}\geq 1-\alpha_{\min}^{2}$. Then $\mathbf{v}_{t,i}\in\overline{S}$ for all $i\geq 2$.
Proof.
Since all $X_{\ell}\in S$ and $S$ is a convex set, $\mathbf{v}_{t,1}\in S$. Now take any $\mathbf{x}\in S$ and $i\geq 2$, and note that
$(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}\leq\sum_{i=2}^{d}(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}=1-(\mathbf{x}^{\top}\mathbf{v}_{t,1})^{2}\leq 1-\alpha_{\min}^{2}\,.$
We use that $\sum_{i=1}^{d}(\mathbf{x}^{\top}\mathbf{v}_{t,i})^{2}=1$ and $\mathbf{x}^{\top}\mathbf{v}_{t,1}\geq\alpha_{\min}$. Therefore, when $\beta_{\max}\geq 1-\alpha_{\min}^{2}$, we have $\mathbf{v}_{t,i}\in\overline{S}$ for all $i\geq 2$. ∎
Our analysis has two parts. First, we bound the approximation error under the assumption that $X_{\ell}\in S$ holds for all $\ell\in[T]$. Second, we show how to choose $\alpha_{\min}$ and $\beta_{\max}$ to guarantee this. We start with the approximation error.
Lemma 8.
Let $X_{\ell}\in S$ hold for all $\ell\in[T]$. Then for any $\mathbf{x}_{*,k}$
$\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{T+1}^{-1}\mathbf{x}_{*,k}\leq\frac{1}{\alpha_{\min}^{2}T+1}+(1-\alpha_{\min}^{2})\,.$
Proof.
We start with
$\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{T+1}^{-1}\mathbf{x}_{*,k}=\sum_{i=1}^{d}\lambda_{T+1,i}^{-1}\mathbf{x}_{*,k}^{\top}\mathbf{v}_{T+1,i}\mathbf{v}_{T+1,i}^{\top}\mathbf{x}_{*,k}\leq\frac{(\mathbf{x}_{*,k}^{\top}\mathbf{v}_{T+1,1})^{2}}{\alpha^{2}_{\min}T+1}+\sum_{i=2}^{d}(\mathbf{x}_{*,k}^{\top}\mathbf{v}_{T+1,i})^{2}\,.$
The inequality uses lower bounds in Lemma 6. Then we apply $\mathbf{x}_{*,k}^{\top}\mathbf{v}_{T+1,1}\leq 1$ and $\sum_{i=2}^{d}(\mathbf{x}_{*,k}^{\top}\mathbf{v}_{T+1,i})^{2}\leq 1-\alpha^{2}_{\min}$. ∎
Now we prove by induction that $X_{\ell}\in S$ holds for all $\ell\in[T]$.
Lemma 9.
Let $X_{\ell}\in S$ hold for all $\ell\in[t-1]$. Suppose that $t\leq\frac{\alpha^{2}_{\min}}{(\beta+\sqrt{2})\beta d}$. Then $\mathbf{x}_{t}\in S$.
Proof.
Our algorithm chooses a example $\mathbf{x}_{t}\in S$ when for all $\mathbf{x}_{*,k}$
$\frac{\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*,k}}{1+\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}}\geq\frac{\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*,k}}{1+\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}}$
holds for any $\mathbf{x}\in S$ and $\mathbf{y}\in\overline{S}$. Since $0\leq\mathbf{v}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{v}\leq 1$ for any $\|\mathbf{v}\|_{2}\leq 1$, the above event occurs when
$\min_{k}(\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x})^{2}=\min_{k}\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}\mathbf{x}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*,k}\geq 2\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}\mathbf{y}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}_{*,k}=2\max_{k\in[K]}(\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y})^{2}\,.$
We start with an upper bound on the right-hand side,
$\max_{k\in[K]}|\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{y}|\leq\max_{k\in[K]}\sum_{i=1}^{d}\lambda_{t,i}^{-1}|\mathbf{x}_{*,k}^{\top}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\mathbf{y}|\leq\beta_{\max}d\,.$
Here we use $\lambda_{t,i}\geq 1$ (Lemma 6), and that $\mathbf{v}_{t,1}^{\top}\mathbf{y}\leq\beta_{\max}$ and $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\leq\beta_{\max}$ when $i\geq 2$.
Now we bound the left-hand side as
$\min_{k}|\mathbf{x}_{*,k}^{\top}\mathbf{\Lambda}_{t}^{-1}\mathbf{x}|\geq\min_{k}\lambda_{t,1}^{-1}|\mathbf{x}_{*,k}^{\top}\mathbf{v}_{t,1}\mathbf{v}_{t,1}^{\top}\mathbf{x}|-\sum_{i=2}^{d}\lambda_{t,i}^{-1}|\mathbf{x}_{*,k}^{\top}\mathbf{v}_{t,i}\mathbf{v}_{t,i}^{\top}\mathbf{x}|$ $\geq\frac{\alpha^{2}_{\min}}{t}-\beta^{2}_{\max}d\,.$
To bound the first term, we use $\lambda_{t,1}\leq t$, and that $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,1}\geq\alpha_{\min}$ and $\mathbf{v}_{t,1}^{\top}\mathbf{x}\geq\alpha_{\min}$. To bound the second term, we use $\lambda_{t,i}\geq 1$, and that $\mathbf{x}_{*}^{\top}\mathbf{v}_{t,i}\leq\beta_{\max}$ and $\mathbf{v}_{t,i}^{\top}\mathbf{x}\leq\beta_{\max}$.
Now we chain all inequalities and get that our algorithm chooses a example $\mathbf{x}_{t}\in S$ when
$t\leq\frac{\alpha^{2}_{\min}}{(\beta_{\max}+\sqrt{2})\beta_{\max}d}\,.$
This completes the proof. ∎
B.4 Proof of Theorem 2
Consider the test examples $\{\mathbf{x}_{*}\}_{k=1}^{K}$. Recall that $H_{t}=\left(X_{\ell},Y_{\ell}\right)_{\ell\in[t-1]}$ is the history till round $t$. Then the posterior variance is
$\widehat{\mathbf{\Sigma}}_{t}=\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}\right)^{-1}$
and $\widehat{\bm{\theta}}_{t}=\widehat{\mathbf{\Sigma}}_{t}\left(\mathbf{\Sigma}_{0}^{-1}\bm{\theta}_{0}+\sigma^{-2}\sum_{\ell=1}^{t-1}X_{\ell}Y_{\ell}\right)$. Now fix a $X_{t}$ at round $t$. Then $Y_{t}=X_{t}^{\top}\bm{\theta}_{*}+\epsilon_{t}$ where $\bm{\theta}_{*}\mid H_{t}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t},\widehat{\mathbf{\Sigma}}_{t}\right)$ and $\epsilon_{t}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. Then $\bm{\theta}_{*}|H_{t+1}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t+1},\widehat{\mathbf{\Sigma}}_{t+1}\right)$ holds for any $Y_{t}$.
Now fix an $\mathbf{x}_{i}\in\mathcal{X}_{\text{examples}}$ and add it to $\widehat{\mathbf{\Sigma}}_{t}$ such that
$\widehat{\mathbf{\Sigma}}_{t,i}=\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\left(\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)\right)^{-1}.$
Define $\hat{\sigma}_{t,i,k}^{2}=\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}^{(j,1)}_{t,i,k}-\tilde{Y}^{(j,2)}_{t,i,k})^{2}$ and the $\mathbb{E}[\hat{\sigma}_{t,i,k}^{2}]=\sigma_{t,i,k}^{2}$. Then define $\sigma_{t,i,\max}^{2}=\max_{k\in[K]}\mathbb{E}[\hat{\sigma}_{t,i,k}^{2}]$. Then using Lemma 10 we can show that with probability $(1-\delta)$
$\sigma_{t,i,\max}^{2}\left[1-2\sqrt{\frac{\log(1/\delta)}{m}}\right]$ $\leq\max_{k\in[K]}\hat{\sigma}_{t,i,k}^{2}\leq\sigma_{t,i,\max}^{2}\left[1+2\sqrt{\frac{\log(1/\delta)}{m}}+\frac{2\log(1/\delta)}{m}\right]$ (4)
Set $m\geq 8\log(1/\delta)$ in (4). It follows then that
$\frac{1}{2}\sigma^{2}_{t,i,\max}\leq\max_{k\in[K]}\frac{1}{m}$ $\sum_{j=1}^{m}(\widetilde{Y}^{(j,1)}_{t,i,k}-\widetilde{Y}^{(j,2)}_{t,i,k})^{2}\leq\frac{5}{2}\sigma^{2}_{t,i,\max}.$
Hence, to minimize the quantity $\max_{k\in[K]}\widetilde{Y}^{(j,1)}_{t,i,k}-\widetilde{Y}^{(j,2)}_{t,i,k}$ we should be minimizing the variance $2\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\sigma^{2}$. Observe that minimizing the variance in SAL leads to minimizing the quantity $2\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\sigma^{2}$ which is same as minimizing the score $\max_{k\in[K]}\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}$ for GO. The claim of the theorem follows.
Lemma 10.
Fix round $t\in[T]$, a sample example $\mathbf{x}_{i}$, and a test example $\mathbf{x}_{*,k}$, and failure probability $\delta\in(0,1)$. Suppose that $m>4\log(1/\delta)$. Define $\hat{\sigma}_{t,i,k}^{2}=\frac{1}{m}\sum_{j=1}^{m}(\tilde{Y}^{(j,1)}_{t,i,k}-\tilde{Y}^{(j,2)}_{t,i,k})^{2}$ and the $\mathbb{E}[\hat{\sigma}_{t,i,k}^{2}]=\sigma_{i,k}^{2}$. Then
$\mathbb{P}\left(\hat{\sigma}_{t,i,k}^{2}\leq\sigma_{i,k}^{2}\left[1-2\sqrt{\frac{\log(1/\delta)}{m}}\right]\right)\leq\delta$
holds with probability at least $1-\delta$. Analogously,
$\mathbb{P}\left(\hat{\sigma}_{t,i,k}^{2}\geq\sigma_{i,k}^{2}\left[1+2\sqrt{\frac{\log(1/\delta)}{m}}+\frac{2\log(1/\delta)}{m}\right]\right)\leq\delta$
holds with probability at least $1-\delta$.
Proof.
Fix an $\mathbf{x}_{i}\in\mathcal{X}_{\text{examples}}$ and add it to $\widehat{\mathbf{\Sigma}}_{t}$. Denote this new co-variance matrix as $\widehat{\mathbf{\Sigma}}_{t,i}$ such that
$\widehat{\mathbf{\Sigma}}_{t,i}=\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\left(\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}+\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right)\right)^{-1}.$
Let $\widetilde{Y}^{(1)}_{t,i,j}=\mathbf{x}_{*,k}^{\top}\bm{\theta}_{*}+\epsilon_{t,i,j,1}$, where $\bm{\theta}_{*}\mid H_{t}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t},\widehat{\mathbf{\Sigma}}_{t+1}\right)$ and $\epsilon_{t,i,j,1}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. This yields that $\widetilde{Y}^{(1)}_{t,i,j}\sim\mathcal{N}\left(\mathbf{x}_{*,k}^{\top}\widehat{\bm{\theta}}_{t+1},\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\sigma^{2}\right)$. Similarly $\widetilde{Y}^{(2)}_{t,i,j}=\mathbf{x}_{*,k}^{\top}\bm{\theta}_{*}+\epsilon_{t,i,j,2}$, where $\bm{\theta}_{*}\mid H_{t}\sim\mathcal{N}\left(\widehat{\bm{\theta}}_{t+1},\widehat{\mathbf{\Sigma}}_{t,i}\right)$ and $\epsilon_{t,i,j,2}\sim\mathcal{N}\left(0,\sigma^{2}\right)$. Therefore, we get that
$\widetilde{Y}^{(j,1)}_{t,i,k}-\widetilde{Y}^{(j,2)}_{t,i,k}\sim\mathcal{N}(0,2\mathbf{x}_{*,k}^{\top}\widehat{\mathbf{\Sigma}}_{t,i}\mathbf{x}_{*,k}+\sigma^{2}).$
Now we proceed the same way as in Lemma 2 of Lalitha et al. [2023]. Using Cochran’s theorem, we have that $\hat{\sigma}_{t,i,k}^{2}m/\sigma_{i,k}^{2}$ is a $\chi^{2}$ random variable with $m$ degrees of freedom. Then using (4.4) and Lemma 1 of Laurent and Massart [2000] we can show that
$\mathbb{P}\left(m-\frac{\hat{\sigma}_{t,i,k}^{2}m}{\sigma_{i,k}^{2}}\geq 2\sqrt{m\log(1/\delta)}\right)\leq\delta$ (5)
Dividing both sides of (5) in the probability by $m$, and multiplying by $\sigma_{i,k}^{2}$, we can get the following
$\mathbb{P}\left(\sigma_{i,k}^{2}(1-2\sqrt{\log(1/\delta)/m})\geq\hat{\sigma}_{t,i,k}^{2}\right)\leq\delta.$
Observe that $01-2\sqrt{\log(1/\delta)/m}>$ 0, we can divide both sides by it and get the first claim of the lemma. The second claim is proved by (4.3) in Laurent and Massart [2000], an immediate corollary of their Lemma 1, we have
$\mathbb{P}\left(\frac{\hat{\sigma}_{t,i,k}^{2}m}{\sigma_{i,k}^{2}}-m\geq 2\sqrt{m\log(1/\delta)}+2\log(1/\delta)\right)\leq\delta$
The claim of the lemma follows. ∎
Appendix C Additional Experiments and Results
We use NVIDIA GeForce RTX 3090 GPU with 24GB RAM to load the Large Language Models for inference. The Mistral-7B model requires less than 16GB RAM, and Vinuna-13B model requires less than 22 GB RAM during execution. To run Falcon-40B model we use AWS ml-g5.12xlarge machine. To run the full set of experiments it takes 24-27 hours of compute job. We now briefly discuss the various datasets used in this work.
C.1 Datasets
We now briefly describe the datasets used for our experiments. All the real-life datasets are from UCI [Markelle Kelly, 1988] and OpenML [Vanschoren et al., 2013] repositories. We use $4$ classification and $3$ regression datasets from UCI and OpenML. Additionally, we use $2$ custom datasets for movie names and entity names for classification task in our experiments. These are as follows:
(1) Iris: We use this UCI dataset for classification task. This dataset consists of four features of flowers and three classes of flowers. We use all four features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $150$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
(2) Banknote-authentication: We use this OpenML dataset for classification task. This dataset consists of five features of banknotes and two classes for identifying fake or original banknote. Out of these five features, we use four features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $150$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
(3) Balance-scale: We use this OpenML dataset for classification task. This dataset consists of five features of a scale and three classes of whether the scale tips left/right or is balanced. Out of these five features, we use four features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $625$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
(4) Thyroid-new: We use this OpenML dataset for classification task. This dataset consists of six features for thyroids and three classes. Out of these six features, we use five features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $215$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
(5) Movie-name: We use this custom dataset for classification task. This dataset consists of movie names across five genres (classes) romance, horror, thriller, sport, and action. We convert the movie names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the movie names and the goal is to identify the common genre. The dataset consists of $100$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
(6) Movie-theme: We use this custom dataset for classification tasks in identifying a common theme between pairs of movies. This dataset consists of movie names across five themes (classes) good-vs-evil, man-vs-nature, redemption, Love conquers all, and coming-of-age. We convert the movie names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the pair of movie names and the goal is to identify the common theme. The dataset consists of $100$ instances. We randomly choose $K=20$ as test examples and the remaining instances as training examples.
(7) Entity-name: We use this custom dataset for classification task. This dataset consists of entity names across five entity types (classes) like mountains, seas, rivers, vehicles, and celebrities. Again, we convert the entity names into $768$ dimensional feature embeddings using Instructor embeddings. Note that in the prompt to the LLM we only pass the entity names and the goal is to identify the entity type. The dataset consists of $100$ pairs of instances. We randomly choose $K=20$ pairs as test examples and the remaining instances as training examples.
(8) Fifa: We use this OpenML dataset for the regression task. This dataset consists of six features of players and the clubs they joined as targets. Out of these six features, we use five features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $18063$ instances. We randomly choose $K=20$ test examples and another $200$ examples as training examples.
(9) Machine-cpu: We use this OpenML dataset for regression tasks. This dataset consists of seven features of machine cpu and the target variable is the performance of the cpu. Out of these seven features, we use five features in the prompts as well as estimating the score for selecting the next action. The dataset consists of $209$ instances. We randomly choose $K=20$ test examples and the remaining examples as training examples.
C.1.1 ARC Experiment
In the recent works of Mirchandani et al. [2023], Srivastava et al. [2022] they showed that LLMs behave as general pattern-matching machine. In fact they showed that LLMs can be used to solve tasks from Abstract Reasoning Corpus (ARC) tasks. In the following experiments, we choose two such tasks: (1) ARC expansion and contraction experiment and (2) ARC rotation experiment.
(1) ARC expansion and contraction experiment: In the expansion and contraction experiment, there are two sets of matrices of dimension $4\times 4$ which constitute half the examples of the feature space $\mathcal{X}$. The first set of input matrices have integer values in their center $2\times 2$ cells while all the other cells are 00. The label space $\mathcal{Y}$ of this $4\times 4$ matrix is also a $4\times 4$ matrix where the $4$ inner cells have moved to the $4$ corners. These matrices are termed as expansion matrices.
Similarly, the other set of $4\times 4$ matrices have the $4$ non-zeros values in their corners. These constitute the remaining examples in $\mathcal{X}$. Then the label space $\mathcal{Y}$ is given by $4\times 4$ matrix where the four non-zeros cells come to the center and all the other cell values are 00. These matrices are termed as contraction matrices. This is shown in Figure 1(a).
Therefore, the feature space $\mathcal{X}$ consists of both the expansion and contraction matrices. At every trial, and $n$ training examples and $K$ test examples are chosen randomly from $\mathcal{X}$. Then we run all baselines for $T$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Table 3.
(2) ARC rotation experiment: In the rotation experiment, there are again two sets of matrices of dimension $4\times 4$ which constitute half the examples of the feature space $\mathcal{X}$. The first set of matrices are have integer values in their four corner cells while all the other cells are 00. The label space $\mathcal{Y}$ of this $4\times 4$ matrix is also a $4\times 4$ matrix where the $4$ corner cell values have moved $90^{\circ}$ in the clockwise direction. These matrices are termed as clockwise matrices.
Similarly, the other set of $4\times 4$ matrices have the $4$ non-zeros values in their corners. These constitute the remaining examples in the feature space $\mathcal{X}$. Then the label space $\mathcal{Y}$ is given by $4\times 4$ matrix where the four non-zeros cells have moved $15^{\circ}$ in the anti-clockwise direction and all the other cell values are 00. These matrices are termed as anti-clockwise matrices. This is shown in Figure 1(b).
Therefore, the feature space $\mathcal{X}$ consists of both the clockwise and anti-clockwise matrices. At every trial, $n$ training examples and $K$ test examples are chosen randomly. Then we run all the baselines for $T$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Table 3.
(a) Classification on ARC Expansion and Contraction Experiment
(b) Classification on ARC Rotation Experiment
Figure 1: Explanation of ARC tasks
C.1.2 PCFG Experiment
In this experiment the goal is to predict the next output of a sequence. In the following experiments we choose two such tasks: (1) PCFG add-subtract experiment and (2) PCFG repeat experiment.
(1) PCFG add-subtract experiment: In the add-subtract experiment, there are two sets of sequence of $4$ integers. The first set of sequence of $4$ integers consists of odd integer values which constitute half the examples in the feature space $\mathcal{X}$. The label space $\mathcal{Y}$ of this sequence of $4$ odd examples is sequences of $5$ integers where the last integer is padded to the original sequence by adding one to the last odd integer. These sequences are termed as add examples.
Similarly, the other set of examples consists of a sequence of $4$ even integer values which constitute the remaining examples in the feature space $\mathcal{X}$. The label space $\mathcal{Y}$ of this sequence of $4$ even integer examples is a sequence of $5$ integers where the last integer is padded to the original sequence by subtracting one from the last even integer. These examples are termed as even examples. This is shown in Figure 2(a).
Therefore, the feature space $\mathcal{X}$ consists of both the odd and even sequence of $4$ integer value examples. At every trial, $n$ training and $K$ test examples are chosen randomly. Then we run all the baselines for $T$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Table 3.
(2) PCFG repeat experiment: In the repeat experiment, there are two sets of sequence of $4$ integers. The first set of sequence of $4$ odd integer values constitute half the examples of the feature space $\mathcal{X}$. The label space $\mathcal{Y}$ of this sequence of $4$ integers examples is a sequence of $5$ integers where the last integer is padded to the original sequence by repeating the first odd integer. These sequences are termed as odd-repeat examples.
Similarly, the other set of examples consists of sequence of $4$ even integer values which constitute the remaining examples in the feature space $\mathcal{X}$. The label space $\mathcal{Y}$ of these sequence of $4$ integer value examples is a sequence of $5$ integers where the last integer is padded to the original sequence by repeating the second even integer. These examples are termed as even-repeat examples. This is shown in Figure 2(b).
Therefore, the feature space $\mathcal{X}$ consist of both the odd-repeat and even-repeat examples. At every trial, $n$ training examples and $K$ test examples are chosen randomly. Then we run all the baselines for $T$ iterations where the classification accuracy is calculated if the LLM is able to predict the exact matching. This experiment is shown in Table 3.
(a) Classification on PCFG add-subtract Experiment
(b) Classification on PCFG repeat Experiment
Figure 2: Explanation of PCFG task.
Appendix D Prompt Examples
Classification Dataset Prompts : Below we give an example of how we use the prompts to be used in the LLM for the Iris misclassification task. Similar types of prompts can be found in Dinh et al. [2022], Suzgun et al. [2022]. This is shown in Figure 3(a). Note that since we have the feature representation of the training and test examples from the dataset, we directly use them as $\mathbf{x}_{i}$ and $\mathbf{x}_{*,k}$.
Regression Dataset Prompts : In Figure 3(b) we give an example of a prompt for regression task in Fifa dataset. Note that since we have the feature representation of the training and test examples from the dataset, we directly use them as $\mathbf{x}_{i}$ and $\mathbf{x}_{*,k}$.
(a) Classification Prompt
(b) Regression Prompt
(c) Movie Prompt
(d) Entity Prompt
Figure 3: Prompt examples for Classification, Regression, Movie, and Prompt
(a) Theme Prompt
(b) PCFG Prompt
Figure 4: Prompt examples for Theme and PCFG tasks
(a) ARC Prompt
Figure 5: Prompt examples for ARC task
[…]
Movie Name Experiment: We use a similar technique as in Iris dataset for this setting. The labels of the movie genres belong to $5$ classes as follows: romance, horror, thriller, sport, and action. At every iteration we pass a set of test movie name examples where each $\mathbf{x}_{*,k}$ is now movie name. Note that we feed the natural language text to the LLM as prompts as shown in Figure 3(c). However, to run GO, SAL, and other baselines we require a featurization of these natural language prompts. We obtain a $768$ dimensional featurized representation of the movie names using Instructor embedding [Su et al., 2022]. This constitutes $\mathbf{x}_{i}\in\mathbb{R}^{768}$ and $\mathbf{x}_{*,k}\in\mathbb{R}^{768}$.
Entity Name Experiment: The labels of the entity genres belong to $5$ classes as follows: mountains, seas, rivers, vehicles, and celebrities. At every iteration, we pass a set of test entity name examples where each $\mathbf{x}_{*,k}$ is now an entity name. Note that we feed the natural language text to the LLM as prompts as shown in Figure 3(d). However, to run GO, SAL, and other baselines we require a featurization of these natural language prompts. Again, we obtain a $768$ dimensional featurized representation of the entity names using Instructor embedding [Su et al., 2022]. This constitutes $\mathbf{x}_{i}\in\mathbb{R}^{768}$ and $\mathbf{x}_{*,k}\in\mathbb{R}^{768}$.
PCFG Experiment: We show an example of this prompt in Figure 4(b). Here we concatenate the sequence to obtain training examples $\mathbf{x}_{i}$ and test examples $\mathbf{x}_{*,k}$. So a sequence of $4$ integers of length $4$ will be represented by $\mathbf{x}_{i},\mathbf{x}_{*,k}\in\mathbb{R}^{16}$. Similarly the label $Y_{i}$ and $Y_{*,k}$ consist of sequence of $5$ integers of length $4$ which we concatentate to get a vector of length $\mathbb{R}^{20}$.
ARC Experiment: We show an example of this prompt in Figure 5(a). Here we vectorized the $4\times 4$ matrix to obtain training examples $\mathbf{x}_{i}\in\mathbb{R}^{16}$ and test examples $\mathbf{x}_{*,k}\in\mathbb{R}^{16}$. Similarly the label $Y_{i}$ and $Y_{*,k}$ consist of vectorized matrices of length $\mathbb{R}^{16}$.
Appendix E Table of Notations
Notations Definition
$n$ Total unlabeled examples
$d$ Dimension of the feature
$\mathcal{X}$ Feature set
$\mathcal{Y}$ Label space
$\bm{\theta}_{*}$ Unknown model parameter
$\mathbf{x}_{i}$ Feature of sample example $i$
$\mathbf{x}_{*,k}$ $k$-th test example
$f(\mathbf{x},\bm{\theta}_{*})$ Model
$Y_{*}$ Label
$H_{t}=(X_{\ell},Y_{\ell})_{\ell\in[t-1]}$ History of $t-1$ previously labeled examples
$p(\cdot\mid\mathbf{x},H_{t})$ Distribution of the label of example $\mathbf{x}$ conditioned on $H_{t}$
$\bm{\theta}_{0}$ Prior mean of the unknown model parameter $\bm{\theta}_{*}$
$\mathbf{\Sigma}_{0}$ Prior mean of the unknown model parameter $\bm{\theta}_{*}$
$\widehat{\mathbf{\Sigma}}_{t}=\left(\mathbf{\Sigma}_{0}^{-1}+\sigma^{-2}\sum_{\ell=1}^{t-1}X_{\ell}X_{\ell}^{\top}\right)^{-1}$ Posterior covariance
$\mathcal{L}_{t}$ Set of labeled examples
$\mathcal{U}_{t}$ Set of unlabeled examples
$\widehat{\bm{\theta}}_{t,i,j}$ Posterior mean
Table 5: Table of Notations